### v1

# Compositional Design of Safety Controllers for Large-scale Stochastic Hybrid Systems

Abstract.
In this work, we propose a compositional scheme based on small-gain reasoning for the safety controller synthesis of interconnected stochastic hybrid systems with both continuous evolutions and instantaneous jumps. In our proposed setting, we first offer an augmented scheme to represent each stochastic hybrid subsystem with continuous and discrete evolutions in a unified framework, ensuring that the state trajectories match those of the original hybrid systems. We then introduce the concept of augmented control sub-barrier certificates (A-CSBC) for each subsystem, which allows the construction of augmented control barrier certificates (A-CBC) for interconnected systems and their safety controllers under small-gain compositional conditions. We eventually leverage the constructed A-CBC and quantify a guaranteed probabilistic bound across the safety of the interconnected system. While the computational complexity of designing a barrier certificate and its safety controller grows polynomially with network dimension using sum-of-squares (SOS) optimization program, our compositional approach significantly reduces it to a linear scale with respect to the number of subsystems. We verify the efficacy of our proposed approach over an interconnected stochastic hybrid system composed of $1000$ nonlinear subsystems.
1. Introduction
Stochastic hybrid systems (SHS) [BLE+06, CL06] are complex heterogeneous models, spanning a wide range of safety-critical applications, in which discrete dynamics describe computation and communication parts and continuous models capture physical components. Providing a formal verification and controller synthesis framework for this type of complex heterogeneous systems is naturally very challenging primarily due to (i) the stochastic nature of dynamics, (ii) the tight interaction between continuous and discrete components (in both space and time), (iii) the large dimension of state/input sets, and (iv) the need to manage complex logic requirements [BK08].
Discretization-based approaches. In order to deal with the above-mentioned difficulties, one potential solution, proposed in the relevant literature, is to approximate original models by a simpler one with either a lower dimension (a.k.a., infinite abstractions) or with discrete-state sets (a.k.a., finite abstractions) [JP09, APLS08]. On the downside, the existing abstraction-based techniques rely on discretizing state/input sets and are not applicable to real-world high-dimensional systems due to the state-explosion problem. To make the abstraction-based techniques more scalable, compositional approaches have been proposed in the relevant literature that construct abstractions of large-scale systems based on those of smaller subsystems [LZ22, HHHK13, NZ20, LSAZ22, Lav24].
Discretization-free approaches. To mitigate the computational complexity arising from abstraction-based approaches, control barrier certificates (CBC), a promising discretization-free approach, have garnered significant attention in recent years for the formal analysis of complex dynamical systems. Specifically, barrier certificates are Lyapunov-like functions whose initial level set separates an unsafe region from all system trajectories starting from a set of initial states. Consequently, the existence of such a function provides a (probabilistic) safety certificate over the system (cf. Fig. 1). This approach has been initially proposed by [PJ04, PJP07] and then widely employed for formal verification and synthesis of non-stochastic [BWAE15, ACE+19] and stochastic dynamical systems [ZRHH10, ASBA19, SDC19, Cla19, NZ22, LF24, NPNS24, ALZ22, JLZ22, WHL24], to name a few.
Figure 1. A control barrier certificate for stochastic dynamical systems. The initial level set of CBC is denoted by dashed line (i.e., $\mathcal{B}(x)=\mu$).
Innovative findings. Despite the extensive above-mentioned literature on barrier certificate construction, there are no results available for large-scale stochastic hybrid systems. Our primary contribution is to propose a compositional scheme for the safety controller synthesis of interconnected SHS with both continuous evolutions, governed by stochastic differential equations with Brownian motions and Poisson processes, and instantaneous jumps, modeled by stochastic difference equations with additive noises. In our proposed setting, we first introduce an augmented framework to recast each stochastic hybrid subsystem with continuous and discrete evolutions via a unified framework, spanning both scenarios. We then introduce notions of augmented control sub-barrier certificates, computed for each subsystem, using which one can construct augmented control barrier certificates for interconnected systems together with their safety controllers under some small-gain compositional conditions. We eventually leverage the constructed augmented control barrier certificate and offer a certified probabilistic bound across the safety of the interconnected system. We verify the usefulness of our proposed approaches by applying them to an interconnected stochastic hybrid system consisting of $1000$ nonlinear subsystems. While the primary results of this work were initially presented in [LSF22], they are limited to monolithic systems and cannot handle models with more than six state dimensions due to the computational complexity involved in searching for control barrier certificates. In contrast, we propose here an innovative compositional technique to address large-scale SHS by constructing an augmented control barrier certificate for the interconnected system based on certificates of individual subsystems (cf. the $1000$-dimensional case study). While the computational complexity of designing a barrier certificate and its safety controller in [LSF22] grows polynomially with system dimension using sum-of-squares (SOS) optimization program, our compositional approach here significantly reduces it to a linear scale with respect to the number of subsystems. Additionally, we present the proofs for all statements that were not included in [LSF22].
Organization. The rest of the paper is structured as follows. Section 2 provides mathematical preliminaries and notations, and also the formal definition of (augmented) stochastic hybrid systems. In Section 3, we introduce notions of augmented control sub-barrier and barrier certificates for, respectively, augmented stochastic hybrid systems with and without disturbance signals. In Section 4, we propose a compositional approach based on small-gain reasoning to construct an augmented control barrier certificate for the interconnected system based on the certificates of individual subsystems. In Section 5, we provide required conditions for the construction of augmented control sub-barrier certificates based on control barrier certificates of original SHS. In Section 6, we employ sum-of-squares (SOS) optimization techniques to search for control barrier certificates of SHS. Section 7 verifies the proposed approaches over a nonlinear interconnected SHS, and concluding remarks are finally provided in Section 8.
2. Stochastic Hybrid Systems
2.1. Notation
Sets of non-negative and positive integers are, respectively, denoted by :=$\mathbb{N}:=\{0,1,2,...\}$ and +=$\mathbb{N}^{+}=\{1,2,...\}$. Sets of real, positive and non-negative real numbers are represented by $\mathbb{R},\mathbb{R}^{+}$, and $\mathbb{R}_{0}^{+}$, respectively. Given $N$ vectors $x_{i}\in\mathbb{R}^{n_{i}}$, $x=[x_{1};...;x_{N}]$ represents the corresponding vector of dimension $\sum_{i}n_{i}$. Given a vector $x\in\mathbb{R}^{n}$ , we define the Hadamard power as $x^{\circ m}=[x_{1}^{m}~{}...~{}x_{n}^{m}]^{\top}$ for $m\in\mathbb{N}$. A (block) diagonal matrix in ${\mathbb{R}}^{N\times{N}}$ with diagonal matrix entries $(A_{1},\ldots,A_{N})$ and scalar entries $(a_{1},\ldots,a_{N})$ is denoted by $\mathsf{blkdiag}(A_{1},\ldots,A_{N})$ and $\mathsf{diag}(a_{1},\ldots,a_{N})$, respectively. We denote the Euclidean norm of a vector $x\in\mathbb{R}^{n}$ by $\|x\|$. Given sets $X_{i},i\in\{1,\ldots,N\}$, their Cartesian product is denoted by $\prod_{i=1}^{N}X_{i}$. Symbol $\mathds{I}_{n}$ denotes the identity matrix in $\mathbb{R}^{n\times{n}}$. Given a measurable function $f:\mathbb{N}\rightarrow\mathbb{R}^{n}$, the (essential) supremum of $f$ is denoted by $\|f\|_{\infty}:=\text{(ess)sup}\{\|f(k)\|,k\geq 0\}$.
2.2. Preliminaries
Given a probability space $(\Omega,\mathcal{F}_{\Omega},\mathds{P}_{\Omega})$, with $\Omega$ being the sample space, $\mathcal{F}_{\Omega}$ a sigma-algebra on $\Omega$, and $\mathds{P}_{\Omega}$ a probability measure, we assume that $(\Omega,\mathcal{F}_{\Omega},\mathds{P}_{\Omega})$ is endowed with a filtration $\mathds{F}=(\mathcal{F}_{s})_{s\geq 0}$ fulfilling conditions of completeness and right continuity. We consider $(\mathbb{W}_{s})_{s\geq 0}$ as a b-dimensional $\mathds{F}$-Brownian motion and $(\mathbb{P}_{s})_{s\geq 0}$ as an r-dimensional $\mathds{F}$-Poisson process. The Poisson process and Brownian motion are considered to be mutually independent. The Poisson process $\mathbb{P}_{s}=[\mathbb{P}_{s}^{1};\cdots;\mathbb{P}_{s}^{\textsf{r}}]$ models r events with independent occurrences.
2.3. Stochastic Hybrid Systems
In this work, our focus lies on stochastic hybrid systems, treating them as individual subsystems, as delineated in the following definition.
Definition 2.1.
A stochastic hybrid system (SHS) $\Phi_{i}$ is characterized by $\Phi_{i}\!=\!(X_{i},U_{i},\mathcal{U}_{i},W_{i},\mathcal{W}_{i},\sigma_{i},\rho_{i},f_{1_{i}},$ $\varsigma_{i},f_{2_{i}})$, where:
  • $X_{i}\subseteq\mathbb{R}^{n_{i}}$ is the state set of SHS;
  • $U_{i}\subseteq\mathbb{R}^{m_{i}}$ is the input set of SHS;
  • $\mathcal{U}_{i}$ is the set of all measurable bounded input functions $\nu_{i}:\mathbb{R}_{0}^{+}\rightarrow U_{i}$;
  • $W_{i}\subseteq\mathbb{R}^{p_{i}}$ is the disturbance set of SHS (which is utilized later for the sake of interconnection);
  • $\mathcal{W}_{i}$ is the set of all measurable bounded disturbance functions $w_{i}:\mathbb{R}_{0}^{+}\rightarrow W_{i}$;
  • $\sigma_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{i}\times\textsf{b}_{i}}$ is the diffusion term which is globally Lipschitz continuous;
  • $\rho_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{i}\times\textsf{r}_{i}}$ is the reset term which is globally Lipschitz continuous;
  • $f_{1_{i}}:X_{i}\times U_{i}\times W_{i}\rightarrow X_{i}$ is the drift term which is globally Lipschitz continuous;
  • $\varsigma_{i}$ is a sequence of independent and identically distributed (i.i.d.) random variables from the sample space $\Omega_{i}$ to the measurable space $(\mathcal{V}_{\varsigma_{i}},\mathcal{F}_{\varsigma_{i}})$:
    $\varsigma_{i}:=\Big{\{}\varsigma_{i}(\cdot)\!:(\Omega_{i},\mathcal{F}_{\Omega_{i}})\rightarrow(\mathcal{V}_{\varsigma_{i}},\mathcal{F}_{\varsigma_{i}})\Big{\}};$
  • $f_{2_{i}}:X_{i}\times U_{i}\times W_{i}\times\mathcal{V}_{\varsigma_{i}}\rightarrow X_{i}$ is the transition map which is globally Lipschitz continuous.
The SHS $\Phi_{i}$ can be characterized by the following stochastic differential and difference equations:
$\Phi_{i}\!:\left\{\begin{array}[]{rl}\mathsf{d}x_{i}(t)\!=\!\!&f_{1_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t))\mathsf{d}t\!+\!\sigma_{i}(x_{i}(t))\mathsf{d}\mathbb{W}_{t}+\rho_{i}(x_{i}(t))\mathsf{d}\mathbb{P}_{t},~{}\quad\quad\quad\quad~{}~{}~{}\!\!\!\!t\in\mathbb{R}_{0}^{+}\backslash\Delta,\\ x_{i}(t)\!=\!\!&f_{2_{i}}(x_{i}(t^{-}),\nu_{i}(t),w_{i}(t),\varsigma_{i}(t)),\hskip 144.54004pt\!\!\!t\in\Delta,\end{array}\right.$ (2.3)
where $\Delta=\{t_{k}\}_{k\in\mathbb{N}}$ with $t_{k+1}-t_{k}\in\{\varepsilon_{1_{i}}\tau,\ldots,\varepsilon_{2_{i}}\tau\}$ for jump parameters $\tau\in\mathbb{R}^{+}$ and $\varepsilon_{1_{i}},\varepsilon_{2_{i}}\in\mathbb{N}^{+}$, $\varepsilon_{1_{i}}\leq\varepsilon_{2_{i}}$. Moreover, $f_{2_{i}}(x_{i}(t^{-}))=\lim_{t\to t_{k}^{-}}f_{2_{i}}(x_{i}(t))$, indicating the value of $f_{2_{i}}(x_{i}(t))$ as $t$ goes to $t_{k}$ from left. Furthermore, $x_{i}:\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}^{n_{i}}$ is the state signal (right-continuous for all $t\in\mathbb{R}_{0}^{+}$) and $\nu_{i}(\cdot)\in\mathcal{U}_{i},w_{i}(\cdot)\in\mathcal{W}_{i}$ are input and disturbance signals. The random sequence $x_{x_{0_{i}}\nu_{i}w_{i}}(t)$ fulfilling (2.3) is called the state trajectory (a.k.a. solution) of $\Phi_{i}$ at time $t\in\mathbb{R}_{0}^{+}$ under signals $\nu_{i}(\cdot),w_{i}(\cdot)$ starting from $x_{0_{i}}$.
Given that control inputs in practice are often fed by a digital controller, we are interested here in sampled-data hybrid systems as
$\mathcal{U}_{\tau_{i}}\!\!=\!\Big{\{}\nu_{i}\!:\mathbb{R}_{0}^{+}\!\rightarrow\!U_{i}\,\big{|}\,\nu_{i}(t)\!=\!\nu_{i}(t_{k-1}),t\!\in\![t_{k-1},t_{k}],k\!\in\!\mathbb{N}\Big{\}},$ (2.4)
with $t_{k}=k\tau$.
2.4. Interconnected SHS
Since the ultimate goal in this work is to synthesize a safety controller for an interconnected SHS, we provide next a formal definition for interconnected SHS without disturbance signals $w$ that can be constructed as a composition of multiple SHS with disturbance signals.
Definition 2.2.
Consider $N\!\in\!\mathbb{N}^{+}$ SHSs $\Phi_{i}=(X_{i},U_{i},$ $\mathcal{U}_{i},W_{i},\mathcal{W}_{i},\sigma_{i},\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$, with their disturbances partitioned as
$w_{i}$ $=[{w_{i1};\ldots;w_{i(i-1)};w_{i(i+1)};\ldots;w_{iN}}].$ (2.5)
The interconnection of $\Phi_{i}$ is $\Phi=(X,U,\mathcal{U},\sigma,\rho,f_{1},\varsigma,f_{2})$, denoted by $\mathcal{I}(\Phi_{1},\ldots,\Phi_{N})$, such that $X=\Pi_{i=1}^{N}X_{i},U=\Pi_{i=1}^{N}U_{i},\sigma:=\mathsf{blkdiag}(\sigma_{1}(x_{1}),\ldots,\sigma_{N}(x_{N})),\rho:=\mathsf{blkdiag}(\rho_{1}(x_{1}),\ldots,\rho_{N}(x_{N})),f_{1}:=[f_{1_{1}};\ldots;f_{1_{N}}],f_{2}:=[f_{2_{1}};\ldots;\\ f_{2_{N}}]$ and $\varsigma:=[\varsigma_{1};\dots;\varsigma_{N}]$, subjected to the following interconnection constraint:
$\forall i,j\in\{1,\dots,N\},i\neq j\!:~{}~{}w_{ij}=x_{j},~{}~{}X_{j}\subseteq W_{ij},$ (2.6)
where $W_{i}:=\prod_{j}W_{ij}$. Such an interconnected SHS can be described by
$\Phi\!:\left\{\begin{array}[]{rl}\mathsf{d}x(t)=\!\!&f_{1}(x(t),\nu(t))\mathsf{d}t+\sigma(x(t))\mathsf{d}\mathbb{W}_{t}+\rho(x(t))\mathsf{d}\mathbb{P}_{t},~{}\quad\quad\quad~{}t\in\mathbb{R}_{0}^{+}\backslash\Delta,\\ x(t)=\!\!&f_{2}(x(t^{-}),\nu(t),\varsigma(t))\hskip 132.87431ptt\in\Delta.\end{array}\right.$ (2.9)
The random sequence $x_{x_{0}\nu}(t)$ fulfilling (2.9) is called the state trajectory of $\Phi$ at time $t\in\mathbb{R}_{0}^{+}$ under an input signal $\nu(\cdot)$ starting from $x_{0}$.
An interconnection of two SHS $\Phi_{i}$ is schematically depicted in Fig. 2.
Figure 2. Interconnection of two SHS $\Phi_{1},\Phi_{2}$.
2.5. Augmented SHS
As it can be observed, the SHS in (2.3) encompasses both continuous and discrete evolutions, in which continuous behaviours are governed by stochastic differential equations with Brownian motions and Poisson processes, and instantaneous jumps are modelled by stochastic difference equations with additive noises. In the next definition, we provide a unified framework for SHS covering both continuous and discrete behaviours by adding an auxiliary variable $\vartheta_{i}$ to the state tuple of $\Phi_{i}$ as a counter, based on which an SHS is allowed or prevented from jump.
Definition 2.3.
Consider an SHS $\Phi_{i}=(X_{i},U_{i},\mathcal{U}_{i},W_{i},\mathcal{W}_{i},\sigma_{i},\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$ with jump parameters ($\tau$, $\varepsilon_{1_{i}}$, $\varepsilon_{2_{i}}$). An augmented SHS (A-SHS) $\mathcal{A}_{i}(\Phi_{i})$ associated to $\Phi_{i}$ is characterized by $\mathcal{A}_{i}(\Phi_{i})\!=\!(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\\ \mathbb{Y}_{i},$ $\mathbb{H}_{i})$, where:
  • $\mathbb{X}_{i}=X_{i}\times\{0,\ldots,\varepsilon_{2_{i}}\}$ is the state space of A-SHS, while $(x_{i},\vartheta_{i})\in\mathbb{X}_{i}$ indicates that the current state is $x_{i}$, and the time elapsed since the latest jump is $\vartheta_{i}$;
  • $\mathbb{U}_{i}=\mathcal{U}_{\tau_{i}}$ is the input space of A-SHS;
  • $\mathds{W}_{i}$ is the disturbance space of A-SHS;
  • $\sigma_{i}$ is the diffusion term;
  • $\rho_{i}$ is the reset term;
  • $\varsigma_{i}$ is a sequence of i.i.d. random variables;
  • $(x_{i}^{\prime},\vartheta_{i}^{\prime})=\mathbb{F}_{i}((x_{i},\vartheta_{i}),\nu_{i},w_{i},\sigma_{i},\rho_{i},\varsigma_{i})$ if and only if one of the following two cases holds:
    1. (i)
      Flow case: $10\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-$ 1, $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, and $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
    2. (ii)
      Jump case: $\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}}$, $x_{i}^{\prime}=f_{2_{i}}(x_{i},\nu_{i},w_{i},\varsigma_{i})$, and $\vartheta_{i}^{\prime}=0$;
  • $\mathbb{Y}_{i}=\mathbb{R}^{n_{i}}$ is the output space of A-SHS;
  • $\mathbb{H}_{i}:\mathbb{X}_{i}\rightarrow\mathbb{Y}_{i}$ is the output map defined as $\mathbb{H}_{i}(x_{i},\vartheta_{i})=x_{i}$.
Remark 2.4.
Note that $(x_{i}^{\prime},\vartheta_{i}^{\prime})=\mathbb{F}_{i}((x_{i},\vartheta_{i}),\nu_{i},w_{i},\sigma_{i},\rho_{i},\varsigma_{i})$ in Definition 2.3 characterizes the evolution of A-SHS, in which $\mathbb{F}_{i}$ is acquired based on one of the following two cases:
- if the counter $\vartheta_{i}$ is between 00 and $\varepsilon_{2_{i}}-1$, then the state evolves based on the flow case as $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, in which $x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$ is the solution of the stochastic differential equation $\mathsf{d}x_{i}(t)=f_{1_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t))\mathsf{d}t\!+\!\sigma_{i}(x_{i}(t))\mathsf{d}\mathbb{W}_{t}+\rho_{i}(x_{i}(t))\mathsf{d}\mathbb{P}_{t},$ which is well defined in the literature [ØS05]. In this case, the next counter is increased as $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
- if the counter $\vartheta_{i}$ is between $\varepsilon_{1_{i}}$ and $\varepsilon_{2_{i}}$, then the state evolves based on the jump case as $x_{i}^{\prime}=f_{2_{i}}(x_{i}(t^{-}),\nu_{i}(t),\\ w_{i}(t),\varsigma_{i}(t))$ and the counter is reset from zero as $\vartheta_{i}^{\prime}=0$.
In the following proposition, we demonstrate that the state trajectories of an A-SHS and its original SHS are equivalent. Consequently, we leverage the A-SHS as a unified framework that encompasses both continuous evolutions and instantaneous jumps, providing a more manageable model.
Proposition 2.5.
The state trajectory of $\Phi_{i}$ defined in (2.3) can be uniquely mapped to a state trajectory of $\mathcal{A}_{i}(\Phi_{i})$ in Definition 2.3 and vice versa.
[…]
3. Augmented Control (Sub-)Barrier Certificates
In this section, we introduce notions of augmented control sub-barrier and barrier certificates for, respectively, augmented SHS with and without disturbance signals. We then utilize those notions and quantify a guaranteed probabilistic bound over the safety of the interconnected SHS.
Definition 3.1.
Given an A-SHS $\mathcal{A}_{i}(\Phi_{i})=(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},\mathbb{H}_{i})$, consider $\mathbb{X}_{0_{i}}=X_{0_{i}}\times\{0\},~{}\mathbb{X}_{u_{i}}=X_{u_{i}}\times\{0,\ldots,\varepsilon_{2_{i}}\}$, as initial and unsafe regions of $\mathcal{A}_{i}(\Phi_{i})$, respectively, with $X_{0_{i}},X_{u_{i}}\subseteq X_{i}$ being initial and unsafe regions of $\Phi_{i}$. A function $\mathcal{B}_{i}:\mathbb{X}_{i}\to{\mathbb{R}}_{0}^{+}$ is said to be an augmented control sub-barrier certificate (A-CSBC) for $\mathcal{A}_{i}(\Phi_{i})$ if there exist $10<\gamma_{i}<$ 1, $\kappa_{i}\in{\mathbb{R}}^{+},\mu_{i},\beta_{i},\varphi_{i},\eta_{i}\in{\mathbb{R}}_{0}^{+}$, such that
$\mathcal{B}_{i}(x_{i},\vartheta_{i})\geq\kappa_{i}\|x_{i}\|^{2},\quad\quad\quad\quad\quad~{}\!\!\forall(x_{i},\vartheta_{i})\in\mathbb{X}_{i},$ (3.1a)
$\mathcal{B}_{i}(x_{i},\vartheta_{i})\leq\mu_{i},\quad\quad\quad\quad\quad\quad\quad\!\!\!\!\forall(x_{i},\vartheta_{i})\in\mathbb{X}_{0_{i}},$ (3.1b)
$\mathcal{B}_{i}(x_{i},\vartheta_{i})\geq\beta_{i},\quad\quad\quad\quad\quad\quad\quad\!\!\!\forall(x_{i},\vartheta_{i})\in\mathbb{X}_{u_{i}},$ (3.1c)
and $\forall(x_{i},\vartheta_{i})\in\mathbb{X}_{i}$, $\exists\nu_{i}\in\mathbb{U}_{i}$, such that $\forall w_{i}=w_{i}(t_{k})\in\mathds{W}_{i}$, $(x_{i}^{\prime},\vartheta_{i}^{\prime})=\mathbb{F}_{i}((x_{i},\vartheta_{i}),\nu_{i},w_{i},\sigma_{i},\rho_{i},\varsigma_{i})$, and
$\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\big{|}x_{i},\nu_{i},w_{i},\vartheta_{i}\Big{]}\!\leq\!\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\varphi_{i}\|w_{i}\|^{2}+\eta_{i},$ (3.1d)
where $\mathds{E}$ is the expected value w.r.t $\varsigma_{i}$ under the one-step transition of $\mathcal{A}_{i}(\Phi_{i})$.
Remark 3.2.
The order of quantifiers used in condition (3.1d), specifically $\forall x_{i}\in X_{i}$, $\exists\nu_{i}\in U_{i}$, $\forall w_{i}\in W_{i}$, inherently facilitates the synthesis of fully-decentralized controllers for $\mathcal{A}_{i}(\Phi_{i})$. This is because the control input $\nu_{i}$ is independent of the internal input $w_{i}$, which represents the state information of other subsystems. As a result, this independence offers considerable flexibility in designing controllers for large-scale networks.
We now provide a notion of augmented control barrier certificates for augmented SHS without disturbance signals, i.e., interconnected A-SHS (cf. Definition 4.1).
Definition 3.3.
Given an interconnected A-SHS $\mathcal{A}(\Phi)=(\mathbb{X},\mathbb{U},\sigma,\rho,\varsigma,\mathbb{F},\mathbb{Y},\mathbb{H})$ with initial and unsafe regions $\mathbb{X}_{0},\mathbb{X}_{u}$, a function $\mathcal{B}:\mathbb{X}\to{\mathbb{R}}_{0}^{+}$ is said to be an augmented control barrier certificate (A-CBC) for $\mathcal{A}(\Phi)$ if there exist $10<\gamma<$ 1, $\mu,\beta,\eta\in{\mathbb{R}}_{0}^{+}$, with $\beta>\mu$, such that
$\mathcal{B}(x,\vartheta)\leq\mu,\quad\quad\quad\quad\quad\quad\quad\quad\!\!\!\forall(x,\vartheta)\in\mathbb{X}_{0},$ (3.2a)
$\mathcal{B}(x,\vartheta)\geq\beta,\quad\quad\quad\quad\quad\quad\quad\quad\!\!\!\forall(x,\vartheta)\in\mathbb{X}_{u},$ (3.2b)
and $\forall(x,\vartheta)\in\mathbb{X}$, $\exists\nu\in U$, such that $(x^{\prime},\vartheta^{\prime})=\mathbb{F}((x,\vartheta),\nu,\sigma,\rho,\varsigma)$, and
$\mathds{E}\Big{[}\mathcal{B}(x^{\prime},\vartheta^{\prime})\,\big{|}\,x,\nu,\vartheta\Big{]}\leq\gamma\mathcal{B}(x,\vartheta)+\eta.$ (3.2c)
Remark 3.4.
It is important to note that merely satisfying conditions (3.1)does not guarantee the safety of the network. In particular, an A-CSBC captures the influence of disturbance inputs $w_{i}$ through interaction gains $\varphi_{i}$ in  (3.1d), reflecting the impact of neighboring subsystems within the network topology. Under certain small-gain compositional conditions detailed in Section 4, these A-CSBCs are combined to form an A-CBC for the interconnected A-SHS, ensuring its safety. To do so, condition (3.1a) is crucial in deriving these compositional results (see (4.3)).
Note that $\beta>\mu$ is required in Definition 3.3 in order to ensure there is no intersection between initial and unsafe sets of interconnected A-SHS. The next theorem, borrowed from [Kus67] but adapted to A-SHS, shows the usefulness of A-CBC by quantifying a guaranteed probabilistic bound on the safety of the interconnected SHS.
Theorem 3.5.
Given interconnected A-SHS $\mathcal{A}(\Phi)=(\mathbb{X},\mathbb{U},\sigma,\rho,\varsigma,\mathbb{F},\mathbb{Y},\mathbb{H})$, let $\mathcal{B}$ be an A-CBC for $\mathcal{A}(\Phi)$. Then for any $x_{0}$ and $\vartheta_{0}$, the probability that the state trajectories of $\mathcal{A}(\Phi)$ reach the unsafe region $\mathbb{X}_{u}$ within the finite-time horizon $k\in[0,\mathcal{T}]$ is computed as:
$\mathds{P}\Big{\{}x_{x_{0}\nu}(t_{k})\in\mathbb{X}_{u}~{}\text{for some}~{}k\in[0,\mathcal{T}]\,\,\big{|}\,\,x_{0},\vartheta_{0}\Big{\}}\leq\begin{cases}1-(1-\frac{\mu}{\beta})(1-\frac{\eta}{\beta})^{\mathcal{T}}\!,\quad\quad&\text{if }\beta\geq\frac{\eta}{1-{\gamma}},\\ (\frac{\mu}{\beta})\gamma^{\mathcal{T}}+(\frac{\eta}{(1-{\gamma})\beta})(1-\gamma^{\mathcal{T}}),\quad\quad&\text{if }\beta<\frac{\eta}{1-{\gamma}}.\\ \end{cases}$ (3.3)
If $\eta=0$ in (3.2c) (i.e., $\mathcal{B}$ being a nonnegative supermartingale), the results of (3.3) can be extended to infinite time horizons as
$\mathds{P}\Big{\{}x_{x_{0}\nu}(t_{k})\in\mathbb{X}_{u}~{}\text{for some}~{}k\in[0,\infty)\,\,\big{|}\,\,x_{0},\vartheta_{0}\Big{\}}\leq\frac{\mu}{\beta}.$ (3.4)
Proof. Based on condition (3.2b), we have $\mathbb{X}_{u}\subseteq\{(x,\vartheta)\in\mathbb{X}\,\,\big{|}\,\,\mathcal{B}(x,\vartheta)\geq\beta\}$. Then one has
$\mathds{P}$$\Big{\{}(x(t_{k}),\vartheta(t_{k}))\in\mathbb{X}_{u}\text{ for some}~{}0\leq k\leq\mathcal{T}\,\,\big{|}\,\,x_{0},\vartheta_{0}\Big{\}}\leq\mathds{P}\Big{\{}\sup_{0\leq k\leq\mathcal{T}}\mathcal{B}(x(t_{k}),\vartheta(t_{k}))\geq\beta\,\,\big{|}\,\,x_{0},\vartheta_{0}\Big{\}}.$ (3.5)
[…]
Searching for control barrier certificates for large-scale dynamical systems is in general computationally very expensive. This has been the main motivation in this work to propose A-CSBC for each $\mathcal{A}_{i}(\Phi_{i})$ instead of searching for A-CBC over the interconnected A-SHS, which is not possible in practice. In the next section, we propose a compositional approach based on small-gain reasoning to construct an A-CBC of the interconnected A-SHS based on A-CSBC of individual subsystems.
4. Compositional Construction of A-CBC
In this section, we analyze networks of A-SHS via small-gain reasoning and discuss how to construct an A-CBC of the interconnected A-SHS based on the corresponding A-CSBCs of individual subsystems. To do so, we first introduce a formal definition for the interconnected A-SHS similar to Definition 2.2.
Definition 4.1.
Given $N\in\mathbb{N}^{+}$ A-SHSs $\mathcal{A}_{i}(\Phi_{i})=(\mathbb{X}_{i},$ $\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},\mathbb{H}_{i})$ with their disturbances partitioned similar to (2.5), the interconnection of $\mathcal{A}_{i}(\Phi_{i})$, $\forall i\in\{1,\ldots,N\}$, is $\mathcal{A}(\Phi)=(\mathbb{X},\mathbb{U},\sigma,\rho,\varsigma,\mathbb{F},\mathbb{Y},\mathbb{H})$, denoted by $\mathcal{I}(\mathcal{A}_{1}(\Phi_{1}),\ldots,\mathcal{A}_{N}(\Phi_{N}))$, such that $\mathbb{X}\!:=\!\prod_{i=1}^{N}\mathbb{X}_{i},\mathbb{U}=\Pi_{i=1}^{N}\mathbb{U}_{i},\sigma\!:=\!\mathsf{blkdiag}(\sigma_{1}(x_{1}),\ldots,\sigma_{N}(x_{N})),$ $\rho:=\mathsf{blkdiag}(\rho_{1}(x_{1}),\ldots,\rho_{N}(x_{N})),\varsigma:=[\varsigma_{1};\dots;\varsigma_{N}],\mathbb{Y}:=\prod_{i=1}^{N}\mathbb{Y}_{i},\mathbb{H}=\prod_{i=1}^{N}\mathbb{H}_{i}$, and the map $\mathbb{F}=[\mathbb{F}_{1};\dots;\mathbb{F}_{N}]$ is defined by $(x^{\prime},\vartheta^{\prime})\!=\!\mathbb{F}((x,\vartheta),\nu,\sigma,\rho,\varsigma)$ if and only if, for any $i\in\{1,\dots,N\}$:
  1. (i)
    Flow case: $10\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-$ 1, $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, and $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
  2. (ii)
    Jump case: $\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}}$, $x_{i}^{\prime}=f_{2_{i}}(x_{i},\nu_{i},w_{i},\varsigma_{i})$, and $\vartheta_{i}^{\prime}=0$;
where $x=[x_{1};\dots;x_{N}],\vartheta=[\vartheta_{1};\dots;\vartheta_{N}],\varepsilon_{1}=[\varepsilon_{1_{i}};\dots;\varepsilon_{1_{N}}],\varepsilon_{2}=[\varepsilon_{2_{i}};\dots;\varepsilon_{2_{N}}]$ and subjected to the following interconnection constraint:
$\forall i,j\in\{1,\dots,N\},i\neq j\!:~{}~{}w_{ij}=x_{j},~{}~{}\mathbb{X}_{j}\subseteq\mathds{W}_{ij},$
where $\mathds{W}_{i}:=\prod_{j}\mathds{W}_{ij}$.
In the next theorem, we propose the main compositionality result of the work, using which one can compositionally construct an A-CBC for the interconnected A-SHS based on A-CSBCs of individual subsystems. To do so, we first define $\Theta:=\mathsf{diag}(\theta_{1},\ldots,\theta_{N})$ with $\theta_{i}=1-\gamma_{i}$, and $\Psi:=\{\psi_{ij}\}$ with $\psi_{ij}=\frac{\varphi_{i}}{\kappa_{j}}$, where $\psi_{ii}=0$, $\forall i\in\{1,\cdots,N\}$.
Theorem 4.2.
Consider an interconnected A-SHS $\mathcal{I}(\mathcal{A}_{1}(\Phi_{1}),\ldots,\mathcal{A}_{N}(\Phi_{N}))$ induced by $N\in\mathbb{N}^{+}$ A-SHSs $\mathcal{A}_{i}(\Phi_{i})$. Suppose that each $\mathcal{A}_{i}(\Phi_{i})$ admits an A-CSBC $\mathcal{B}_{i}$, as defined in Definition 3.1. For a vector $\xi$ with $\xi_{i}>0,\forall i\in\{1,\dots,N\}$, if
$\xi^{\top}(-\Theta+\Psi)$ $:=\big{[}\xi_{1}\Gamma_{1};\dots;\xi_{N}\Gamma_{N}\big{]}^{\top}<0,$ (4.1a)
$\sum_{i=1}^{N}\xi_{i}\beta_{i}$ $>\sum_{i=1}^{N}\xi_{i}\mu_{i},$ (4.1b)
then
$\mathcal{B}(x,\vartheta):=\sum_{i=1}^{N}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})$ (4.2)
is an A-CBC for the interconnected A-SHS with the initial and unsafe sets $\mathbb{X}_{0}:=\prod_{i=1}^{N}\mathbb{X}_{0_{i}}$, $\mathbb{X}_{u}:=\prod_{i=1}^{N}\mathbb{X}_{u_{i}}$, with
$\beta:=\sum_{i=1}^{N}\xi_{i}\beta_{i},\quad\mu:=\sum_{i=1}^{N}\xi_{i}\mu_{i},\quad\eta:=\sum_{i=1}^{N}\xi_{i}\eta_{i},\quad\gamma:=1+\Gamma,~{}\text{with}~{}\max_{1\leq i\leq N}\Gamma_{i}<\Gamma<0,~{}\text{and}~{}\Gamma\in(-1,0).$
Proof. We first show that conditions (3.2a)-(3.2b) hold true. For any $(x,\vartheta)\in\mathbb{X}_{0}=\prod_{i=0}^{N}\mathbb{X}_{0_{i}}$ and from (3.1b), one has
$\mathcal{B}(x,\vartheta)=\sum_{i=1}^{N}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\leq\sum_{i=1}^{N}\xi_{i}\mu_{i}=\mu.$
Moreover, for any $(x,\vartheta)\in\mathbb{X}_{u}=\prod_{i=1}^{N}\mathbb{X}_{u_{i}}$ and from (3.1c), we have
$\mathcal{B}(x,\vartheta)=\sum_{i=1}^{N}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\geq\sum_{i=1}^{N}\xi_{i}\beta_{i}=\beta,$
implying that conditions (3.2a)-(3.2b) are satisfied with $\mu=\sum_{i=1}^{N}\xi_{i}\mu_{i}$ and $\beta=\sum_{i=1}^{N}\xi_{i}\beta_{i}$. Note that $\beta>\mu$ according to (4.1b).
$\mathds{E}$ $\Big{[}\mathcal{B}(x^{\prime},\vartheta^{\prime})\,\big{|}\,x,\nu,\vartheta\Big{]}=\mathds{E}\Big{[}\sum_{i=1}^{N}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\,\big{|}\,x,\nu,\vartheta\Big{]}=\sum_{i=1}^{N}\xi_{i}\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i},\vartheta_{i})\,\big{|}\,x,\nu,\vartheta\Big{]}=\sum_{i=1}^{N}\xi_{i}\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i},\vartheta_{i})\,\big{|}\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big{]}$
$\leq\sum_{i=1}^{N}\xi_{i}\big{(}\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\varphi_{i}\|w_{i}\|^{2}+\eta_{i}\big{)}\leq\sum_{i=1}^{N}\xi_{i}\big{(}\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\varphi_{i}\sum_{j=1,i\neq{j}}^{N}\|w_{ij}\|^{2}+\eta_{i}\big{)}$
$=\sum_{i=1}^{N}\xi_{i}\big{(}\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\varphi_{i}\sum_{j=1,i\neq{j}}^{N}\|x_{j}\|^{2}+\eta_{i}\big{)}\leq\sum_{i=1}^{N}\xi_{i}\big{(}\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\!\!\!\!\!\sum_{j=1,i\neq{j}}^{N}\frac{\varphi_{i}}{\kappa_{j}}\mathcal{B}_{j}(x_{j},\vartheta_{j})+\eta_{i}\big{)}$
$=\sum_{i=1}^{N}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\sum_{i=1}^{N}\xi_{i}\big{(}-\theta_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\sum_{j=1,i\neq{j}}^{N}\psi_{ij}\mathcal{B}_{j}(x_{j},\vartheta_{j})+\eta_{i}\big{)}$
$=\mathcal{B}(x,\vartheta)+\xi^{\top}(-\Theta+\Psi)\underbrace{\big{[}\mathcal{B}_{1}(x_{1},\vartheta_{1});\ldots;\mathcal{B}_{N}(x_{N},\vartheta_{N})\big{]}}_{\tilde{\mathcal{B}}(x,\vartheta)}+\sum_{i=1}^{N}\xi_{i}\eta_{i}\leq\underbrace{(1+\Gamma)}_{\gamma}\mathcal{B}(x,\vartheta)+\eta.$ (4.3)
Now we proceed with showing that condition (3.2c) holds, as well. By leveraging conditions (3.1a),(4.1a), one can acquire the chain of inequalities in (4.3). By defining
$\gamma s$ $:=\max\Big{\{}s+\xi^{\top}(-\Theta+\Psi)\tilde{\mathcal{B}}(x,\vartheta)\,\big{|}\,\xi^{\top}\tilde{\mathcal{B}}(x,\vartheta)=s\Big{\}},\quad\eta:=\sum_{i=1}^{N}\xi_{i}\eta_{i},$
with ~$\tilde{\mathcal{B}}(x,\vartheta)=\big{[}\mathcal{B}_{1}(x_{1},\vartheta_{1});\ldots;\mathcal{B}_{N}(x_{N},\vartheta_{N})\big{]}$, condition (3.2c) is also met. The only remaining part is to show that $\gamma=1+\Gamma$ and $10<\gamma<$ 1. Given that $\xi^{\top}(-\Theta+\Psi):=\big{[}\xi_{1}\Gamma_{1};\dots;\xi_{N}\Gamma_{N}\big{]}^{\top}<0$, and $\max_{1\leq i\leq N}\Gamma_{i}<\Gamma<0$ with $\Gamma\in(-1,0)$, we have
$\gamma s$ $=s+\xi^{\top}(-\Theta+\Psi)\tilde{\mathcal{B}}(x,\vartheta)$
$=s+\big{[}\xi_{1}\Gamma_{1};\dots;\xi_{N}\Gamma_{N}\big{]}^{\top}\big{[}\mathcal{B}_{1}(x_{1},\vartheta_{1});\ldots;\mathcal{B}_{N}(x_{N},\vartheta_{N})\big{]}$
$=s+\xi_{1}\Gamma_{1}\mathcal{B}_{1}(x_{1},\vartheta_{1})+\dots+\xi_{N}\Gamma_{N}\mathcal{B}_{N}(x_{N},\vartheta_{N})$
$\leq s+\Gamma\big{(}\xi_{1}\mathcal{B}_{1}(x_{1},\vartheta_{1})+\dots+\xi_{N}\mathcal{B}_{N}(x_{N},\vartheta_{N})\big{)}$
$=s+\Gamma s=(1+\Gamma)s.$
[…]
As demonstrated in Theorem 4.2, under the compositionality conditions of (4.1a)-(4.1b), an A-CBC for the interconnected A-SHS can be compositionally constructed based on A-CSBCs of individual subsystems. In the next section, we provide the necessary conditions for constructing the A-CSBC for the augmented SHS.
5. Construction of A-CSBC
In this section, we establish conditions on the SHS $\Phi_{i}$ that allow for the construction of an A-CSBC for $\mathcal{A}_{i}(\Phi_{i})$. To do so, we first define a notion of control barrier certificates (CBC) over original $\Phi_{i}$.
Definition 5.1.
Given an SHS $\Phi_{i}$ with initial and unsafe sets $X_{0_{i}},X_{u_{i}}\subseteq X_{i}$, a function ¯$\bar{\mathcal{B}}_{i}:X_{i}\rightarrow\mathbb{R}_{0}^{+}$ is called a control barrier certificate (CBC) for $\Phi_{i}$ if there exist $\gamma_{1_{i}}\in{\mathbb{R}},\bar{\kappa}_{i},\gamma_{2_{i}}\in{\mathbb{R}}^{+}$, $\bar{\mu}_{i},\bar{\beta}_{i},\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}},\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\in{\mathbb{R}}_{0}^{+}$, such that
¯$\bar{\mathcal{B}}_{i}(x_{i})\geq\bar{\kappa}_{i}\|x_{i}\|^{2},\quad\quad\quad\quad~{}~{}~{}~{}~{}\forall x_{i}\in X_{i},$ (5.1a)
¯$\bar{\mathcal{B}}_{i}(x_{i})\leq\bar{\mu}_{i},\quad\quad\quad\quad\quad\quad\!\!\forall x_{i}\in X_{0_{i}},$ (5.1b)
¯$\bar{\mathcal{B}}_{i}(x_{i})\geq\bar{\beta}_{i},\quad\quad\quad\quad\quad\quad\!\forall x_{i}\in X_{u_{i}},$ (5.1c)
  • $\forall x_{i}\in X_{i}$, $\exists\nu_{i}\in U_{i}$ such that $\forall w_{i}\in W_{i}$,
¯$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})\leq-\gamma_{1_{i}}\bar{\mathcal{B}}_{i}(x_{i})+\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\bar{\eta}_{1_{i}},$ (5.1d)
where ¯$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})$ is the infinitesimal generator acting on ¯$\bar{\mathcal{B}}_{i}(x_{i})$, defined as [Oks13]
¯$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})$ =¯$=\frac{\partial\bar{\mathcal{B}}_{i}(x_{i})}{\partial x_{i}}\!f_{1_{i}}\!(x_{i},\!\nu_{i},\!w_{i})+\frac{1}{2}\!\textsf{Tr}~{}\!(\sigma_{i}(x_{i})\sigma_{i}(x_{i})\!^{\top}\frac{\partial^{2}{\bar{\mathcal{B}}_{i}(x_{i})}}{\partial x_{i}\partial x_{j}})+\sum_{j=1}^{\textsf{r}_{i}}\lambda_{j_{i}}(\,\bar{\mathcal{B}}_{i}(x_{i}+\rho_{i}(x_{i})\textsf{e}_{j_{i}}^{\textsf{r}_{i}})-\bar{\mathcal{B}}_{i}(x_{i})),$
where $\lambda_{j}$ is the rate of Poisson processes and $\textsf{e}_{j}^{\textsf{r}}$ is an r-dimensional vector with $1$ on the $j$-th entry and 00 elsewhere;
  • $\forall x_{i}\in X_{i}$, $\exists\nu_{i}\in U_{i}$ such that $\forall w_{i}\in W_{i}$,
$\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(f_{2_{i}}(x_{i},\!\nu_{i},\!w_{i})\!)\big{|}x_{i},\!\nu_{i},\!w_{i}\Big{]}\!\!\leq\!\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\!\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}.$ (5.1e)
Before offering the main result of this section, we also raise the following lemma to quantify an upper bound on the evolution of CBC ¯$\bar{\mathcal{B}}_{i}$.
Lemma 5.2.
Given an SHS $\Phi_{i}=(X_{i},U_{i},\mathcal{U}_{i},W_{i},\mathcal{W}_{i},\sigma_{i},$ $\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$, suppose (5.1d) holds. Then for any $x_{i}\in X_{i},\nu_{i}\in\mathcal{U}_{\tau_{i}},w_{i}\in W_{i}$, and for any two consecutive time instances $(t_{k},t_{k+1})$, one has
$\mathds{E}$ [¯$\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}\leq e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k})\!)+e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}}).$ (5.2)
Proof. From Dynkin’s formula [Dyn65], we have:
$\mathds{E}$ [¯$\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}=\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))\!+\!\mathds{E}\Big{[}\int_{t_{k}}^{t_{k+1}}\!\!\!\!\!\!\!\mathcal{L}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))\mathsf{d}t\Big{]}\!.$
From condition (5.1d), one has
$\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$$\leq\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))+\mathds{E}\Big{[}\int_{t_{k}}^{t_{k+1}}\big{(}-\gamma_{1_{i}}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))+\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\bar{\eta}_{1_{i}}\big{)}\mathsf{d}t$
=¯$=\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))\!+\!(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}})-\int_{t_{k}}^{t_{k+1}}\!\!\gamma_{1_{i}}\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))\Big{]}\mathsf{d}t.$
By employing Gronwall inequality [Gro19], we have
$\mathds{E}$ [¯$\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}\leq e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))+e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}}),$
[…]
Under Definition 5.1 and Lemma 5.2, we propose the next theorem to construct an A-CSBC for $\mathcal{A}_{i}(\Phi_{i})$ based on the CBC ¯$\bar{\mathcal{B}}_{i}$.
Theorem 5.3.
Given an SHS $\Phi_{i}=(X_{i},U_{i},\mathcal{U}_{i},W_{i},\mathcal{W}_{i},$ $\sigma_{i},\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$ with its associated A-SHS $\mathcal{A}_{i}(\Phi_{i})=(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},\mathbb{H}_{i})$, suppose ¯$\bar{\mathcal{B}}_{i}$ is a CBC for $\Phi_{i}$. If
$\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\vartheta_{i}<0,\quad\forall\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\},$ (5.3)
then
$\mathcal{B}_{i}(x_{i},\vartheta_{i})=\varpi_{i}~{}\!\bar{\mathcal{B}}_{i}(x_{i})$ (5.4)
is an A-CSBC for $\mathcal{A}_{i}(\Phi_{i})$ with
$\varpi_{i}:=\left\{\begin{array}[]{lr}1,\quad\quad\quad\quad\,~{}~{}~{}~{}\text{if}~{}\gamma_{1_{i}}>0~{}\&~{}0<\gamma_{2_{i}}<1,\\ e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}},\quad\text{if}~{}\gamma_{1_{i}}>0~{}\&~{}\gamma_{2_{i}}\geq 1,\\ \gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}},\quad\quad\quad\quad~{}~{}~{}~{}\!\!\text{if}~{}\gamma_{1_{i}}\leq 0~{}\&~{}0<\gamma_{2_{i}}<1,\end{array}\right.$ (5.8)
for some $10<\alpha_{1_{i}}<$ 1 and $\alpha_{2_{i}}>\varepsilon_{2_{i}}$.
Proof. By employing (5.1a), for any $(x_{i},\vartheta_{i})\in\mathbb{X}_{i}$, one has
$\mathcal{B}_{i}(x_{i},\vartheta_{i})=\varpi_{i}~{}\!\bar{\mathcal{B}}_{i}(x_{i})\geq\varpi_{i}\bar{\kappa}_{i}\|x_{i}\|^{2}.$
Then condition (3.1a) holds with $\kappa_{i}=\varpi_{i}\bar{\kappa}_{i}$. By leveraging (5.1b), for any $(x_{i},\vartheta_{i})\in\mathbb{X}_{0_{i}}$, one has
$\mathcal{B}_{i}(x_{i},\vartheta_{i})=\varpi_{i}~{}\!\bar{\mathcal{B}}_{i}(x_{i})\leq\varpi_{i}~{}\!\bar{\mu}_{i},$
and similarly by utilizing (5.1c) for any $(x_{i},\vartheta_{i})\in\mathbb{X}_{u_{i}}$, one has
$\mathcal{B}_{i}(x_{i},\vartheta_{i})=\varpi_{i}~{}\!\bar{\mathcal{B}}_{i}(x_{i})\geq\varpi_{i}~{}\!\bar{\beta}_{i},$
satisfying conditions (3.1b) and (3.1c) with $\mu_{i}=\varpi_{i}\bar{\mu}_{i}$ and $\beta_{i}=\varpi_{i}\bar{\beta}_{i}$.
Now we proceed with showing condition (3.1d), as well. From (5.2) with $t_{k+1}=\tau,t_{k}=0$, one gets
$\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$ $\leq e^{-\gamma_{1_{i}}\tau}\big{(}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(0))\!+\!\tau(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}})\big{)}$
$=e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})+e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}.$
In order to show that $\mathcal{B}_{i}$ defined in (5.4) fulfills condition (3.1d), we consider the two cases in Definition 2.3 and the three bounds for different values of $\gamma_{1_{i}}$ and $\gamma_{2_{i}}$ in (5.8) as follows:
  • Bound 1 ($\gamma_{1_{i}}>0~{}\&~{}0<\gamma_{2_{i}}<1$):
    1. (i)
      Flow case $(0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1,\vartheta_{i}^{\prime}=\vartheta_{i}+1)$:
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big{|}\,x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big{]}$ $=\!\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\!\nu_{i},\!w_{i}\Big{]}$
      $\leq e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})\!+\!e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}\!+\!e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}\!+\!\|w_{i}\|^{2})+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}.$
    2. (ii)
      Jump case $(\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}},\vartheta_{i}^{\prime}=0)$:
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big{|}\,x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big{]}\!$ $=\!\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\!\nu_{i},\!w_{i}\Big{]}$
      $\leq\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}=\gamma_{2_{i}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}+\bar{\eta}_{2_{i}}.$
      By defining $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau},\gamma_{2_{i}}\}$, $\varphi_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}$ and $\eta_{i}\!=\!\max\{e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}$, we have
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\!\vartheta_{i}^{\prime})\big{|}x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big{]}$ $\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})+\varphi_{i}\|w_{i}\|^{2}\!+\!\eta_{i}.$
  • Bound 2 ($\gamma_{1_{i}}>0~{}\&~{}\gamma_{2_{i}}\geq 1$):
    1. (i)
      Flow case $(0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1,\vartheta_{i}^{\prime}=\vartheta_{i}+1)$:
      $\hskip 56.9055pt\mathds{E}$ $\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big{]}$
      $=e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}^{\prime}}\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$
      $=e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$
      $\leq e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}\big{(}e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})+e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}\big{)}$
      $=e^{-\gamma_{1_{i}}\tau}\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}}e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}}\bar{\mathcal{B}}_{i}(x_{i})+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}+\|w_{i}\|^{2})+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~{}~{}~{}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}.$
    2. (ii)
      Jump case $(\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}},\vartheta_{i}^{\prime}=0)$:
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big{]}$ $=e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}^{\prime}}\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$
      $=\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$
      $\leq\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
      $=\frac{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}}}{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}}}\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}}\gamma_{2_{i}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}.$
      By defining
      $\gamma_{i}$ $=\max\{e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})},e^{-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}}\gamma_{2_{i}}\},\quad\varphi_{i}=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\},$
      and
      $\eta_{i}$ $=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\},$
      we have
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\big{|}x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big{]}$ $\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\varphi_{i}\|w_{i}\|^{2}\!\!+\!\eta_{i}.$
  • Bound 3 ($\gamma_{1_{i}}\leq 0~{}\&~{}0<\gamma_{2_{i}}<1$):
    1. (i)
      Flow case $(0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1,\vartheta_{i}^{\prime}=\vartheta_{i}+1)$:
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big{]}$
      $=\gamma_{2_{i}}^{\frac{\vartheta_{i}^{\prime}}{\alpha_{2_{i}}}}\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$
      $=\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$
      $\leq\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}\Big{(}e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})+e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}\Big{)}$
      $=e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}}\bar{\mathcal{B}}_{i}(x_{i})+\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}\!+\!\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}\!e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}+\|w_{i}\|^{2})+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}.$
    2. (ii)
      Jump case $(\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}},\vartheta_{i}^{\prime}=0)$:
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big{]}$ $=\gamma_{2_{i}}^{\frac{\vartheta_{i}^{\prime}}{\alpha_{2_{i}}}}\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$
      $=\!\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\big{|}x_{i},\!\nu_{i},\!w_{i}\Big{]}\leq\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
      $=\frac{\gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}}}{\gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}}}\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
      $\leq\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}.$
      By defining
      $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}},\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}\},\quad\varphi_{i}=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\},$
      and
      $\eta_{i}=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\},$
      we have
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\!\vartheta_{i}^{\prime})\big{|}x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big{]}\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})+\varphi_{i}\|w_{i}\|^{2}+\eta_{i}.$
To complete the proof, we proceed with showing that $10<\gamma_{i}<$ 1 for all Bounds 1-3. One can readily verify that
$0<\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau},\gamma_{2_{i}}\}<1,$
for Bound 1. In Bound 2, since $10<\alpha_{1_{i}}<$ 1 and $\gamma_{1_{i}}>0$, then we have $e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}<1$. In addition,
$e^{-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}}\gamma_{2_{i}}<1\Leftrightarrow\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}<0.$
Since $\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\vartheta_{i}<0$ for any $\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\}$ and by leveraging the continuity property of real numbers, one can always find some $10<\alpha_{1_{i}}<$ 1 such that $\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}<0$. Hence, $10<\gamma_{i}<$ 1. Similarly, in Bound 3, one has $\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}<1$ since $\alpha_{2_{i}}>\varepsilon_{2_{i}}$ and $10<\gamma_{2_{i}}<$ 1. Additionally,
$e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}<1\Leftrightarrow\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\alpha_{2_{i}}<0.$
[…]
6. Computation of CBC
Here, we employ sum-of-squares (SOS) optimization techniques [Par03] and search for a CBC and its corresponding controller for each $\Phi_{i}$. To do so, we assume that $f_{1_{i}}$ and $f_{2_{i}}$ are polynomial functions. The next lemma provides the SOS formulation.
Lemma 6.1.
Suppose sets $X_{0_{i}},X_{u_{i}},X_{i},U_{i},$ and $W_{i}$ can be described by $X_{0_{i}}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{0_{i}}(x_{i})\geq 0\}$, $X_{u_{i}}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{u_{i}}(x_{i})\geq 0\}$, $X_{i}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{i}(x_{i})\geq 0\}$, $U_{i}=\{U_{i}\subseteq{\mathbb{R}}^{m_{i}}\mid g_{\nu_{i}}(\nu_{i})\geq 0\}$, and $W_{i}=\{W_{i}\subseteq{\mathbb{R}}^{p_{i}}\mid g_{w_{i}}(w_{i})\geq 0\}$. Let there exist an SOS polynomial ¯$\bar{\mathcal{B}}_{i}(x_{i})$, constants $\gamma_{1_{i}}\in{\mathbb{R}},\bar{\kappa}_{i},\gamma_{2_{i}}\in{\mathbb{R}}^{+}$, $\bar{\mu}_{i},\bar{\beta}_{i},\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}},\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\in{\mathbb{R}}_{0}^{+}$, polynomials $l_{\nu_{j_{i}}}(x_{i})$, $\hat{l}_{\nu_{j_{i}}}(x_{i})$ corresponding to the $j^{\text{th}}$ input in $\nu_{i}=(\nu_{1_{i}},\nu_{2_{i}},\ldots,\nu_{m_{i}})\in U_{i}\subseteq{\mathbb{R}}^{m_{i}}$ and $\bar{\nu}_{i}=(\bar{\nu}_{1_{i}},\bar{\nu}_{2_{i}},\ldots,\bar{\nu}_{m_{i}})\in U_{i}\subseteq{\mathbb{R}}^{m_{i}}$, respectively. Suppose there exist vectors of SOS polynomials $l_{0_{i}}(x_{i})$, $l_{u_{i}}(x_{i})$, $\tilde{l}_{i}(x_{i})$, $l_{i}(x_{i},\nu_{i},w_{i}),\hat{l}_{i}(x_{i},\nu_{i},w_{i}),l_{\nu_{i}}(x_{i},\nu_{i},w_{i}),\hat{l}_{\nu_{i}}(x_{i},\nu_{i},w_{i}),l_{w_{i}}(x_{i},\nu_{i},\\ w_{i}),\hat{l}_{w_{i}}(x_{i},\nu_{i},w_{i})$ of appropriate dimensions such that the following expressions are SOS polynomials:
$\bar{\mathcal{B}_{i}}(x_{i})-\tilde{l}_{i}^{\top}(x_{i})g_{i}(x_{i})-\bar{\kappa}_{i}x_{i}^{\top}x_{i},$ (6.1a)
-- $\bar{\mathcal{B}_{i}}(x_{i})-l_{0_{i}}^{\top}(x_{i})g_{0_{i}}(x_{i})+\bar{\mu}_{i},$ (6.1b)
¯$\bar{\mathcal{B}}_{i}(x_{i})\!-\!l_{u_{i}}^{\top}(x_{i})g_{u_{i}}(x_{i})\!-\!\bar{\beta}_{i},$ (6.1c)
-- ¯$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})\!-\!\gamma_{1_{i}}\!(\bar{\mathcal{B}}_{i}(x_{i})\!)\!+\!\bar{\varphi}_{i_{1}}w_{i}^{\top}w_{i}\!+\!\bar{\eta}_{1_{i}}\!-\!\!\!\sum_{j=1}^{m_{i}}(\nu_{j_{i}}\!-\!l_{\nu_{j_{i}}}(x_{i})\!)-l_{i}^{\top}(x_{i},\nu_{i},w_{i})g_{i}(x_{i})$
$-l_{\nu_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{\nu_{i}}(\nu_{i})-l_{w_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{w_{i}}(w_{i}),$ (6.1d)
-- $\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(f_{2_{i}}(x_{i},\nu_{i},w_{i})\!)\big{|}x_{i},\!\nu_{i},\!w_{i}\Big{]}\!+\!\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}w_{i}^{\top}w_{i}+\bar{\eta}_{2_{i}}\!-\!\sum_{j=1}^{m_{i}}(\nu_{j_{i}}\!-\!\hat{l}_{\nu_{j_{i}}}(x_{i})\!)$
$-\hat{l}_{i}^{\top}(x_{i},\nu_{i},w_{i})g_{i}(x_{i})-\hat{l}_{\nu_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{\nu_{i}}(\nu_{i})-\hat{l}_{w_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{w_{i}}(w_{i}).$ (6.1e)
Then ¯$\bar{\mathcal{B}}_{i}(x_{i})$ fulfills conditions (5.1a)-(5.1e) in Definition 5.1. Furthermore, $\nu_{i}=[l_{\nu_{1_{i}}}(x_{i});\dots;l_{\nu_{m_{i}}}(x_{i})]$ and $\hat{\nu}_{i}=[\hat{l}_{\nu_{1_{i}}}(x_{i});\dots;\hat{l}_{\nu_{m_{i}}}(x_{i})]$ are corresponding controllers in flow and jump cases, respectively.
Proof. Since $\bar{\mathcal{B}_{i}}(x_{i})-\tilde{l}_{i}^{\top}(x_{i})g_{i}(x_{i})-\bar{\kappa}_{i}x_{i}^{\top}x_{i}$ is SOS and the term $\tilde{l}_{i}^{\top}(x_{i})g_{i}(x_{i})$ is non-negative over $X_{i}$, the new condition (6.1a) implies condition (5.1a) in Definition 5.1. Similarly, we can show that conditions (6.1b),(6.1c) imply (5.1b),(5.1c), respectively. Now we show that condition (6.1d) implies (5.1d), as well. By selecting inputs $\nu_{j_{i}}=l_{\nu_{j_{i}}}(x_{i})$ and since $l_{i}^{\top}\!(x_{i},\!\nu_{i},\!w_{i})g_{i}(x_{i})$, $l_{\nu_{i}}^{\top}\!(x_{i},\!\nu_{i},\!w_{i})g_{\nu_{i}}(\nu_{i})$, and $l_{w_{i}}^{\top}\!(x_{i},\!\nu_{i},\!w_{i})g_{w_{i}}\!(w_{i})$ are all non-negative, we have
¯$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})\leq-\gamma_{1_{i}}\bar{\mathcal{B}}_{i}(x_{i})+\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\bar{\eta}_{1_{i}}.$
[…]
Remark 6.2.
The computational complexity of SOS is influenced by both the polynomial degree of CBC and the dimensionality of the state variables. Existing research indicates that when the degree of CBC is fixed, the computational demands increase polynomially as the network dimension grows [WTL15]. In contrast, our compositional data-driven approach significantly reduces this complexity, scaling linearly with the number of subsystems.
7. Case Study
Here, we demonstrate the effectiveness of our proposed results by applying them to an interconnected stochastic hybrid system composed of $1000$ nonlinear subsystems. The model of the interconnected SHS $\Phi$ is given as
$\Phi\!:\left\{\begin{array}[]{rl}\mathsf{d}x(t)\!=\!\!&(A_{1}x^{\circ 3}(t)\!+\!B_{1}\nu(t))\mathsf{d}t+E\mathsf{d}\mathbb{W}_{t}+F\mathsf{d}\mathbb{P}_{t},\quad\quad\quad\quad~{}t\in\mathbb{R}_{0}^{+}\backslash\Delta,\\ x(t)\!=\!\!&A_{2}x^{\circ 3}(t^{-})\!+\!B_{2}\nu(t)\!+\!G\varsigma(t),\quad\quad\quad\quad\quad\quad\quad\quad\quad\!\!\!t\in\Delta,\end{array}\right.$ (7.3)
where $A_{1},A_{2}$ are matrices with diagonal elements $\bar{a}_{1_{ii}}=a_{1},\bar{a}_{2_{ii}}=a_{2}$, off-diagonal elements $\bar{a}_{1_{i,i+1}}=\bar{a}_{1_{i+1,i}}=\bar{a}_{1_{1,N}}=\bar{a}_{1_{N,1}}=d_{1}$, $\bar{a}_{2_{i,i+1}}=\bar{a}_{2_{i+1,i}}=\bar{a}_{2_{1,N}}=\bar{a}_{2_{N,1}}=d_{2}$, $i\in\{2,\ldots,N-1\}$, and all other elements are identically zero. Moreover, $B_{1}=\mathsf{diag}(b_{1})$, $B_{2}=\mathsf{diag}(b_{2})$, $E=0.6\mathds{I}_{N},F=0.5\mathds{I}_{N},G=0.5\mathds{I}_{N}$, $x=[x_{1};\ldots;x_{N}]$, $\nu(t)=[\nu_{1}(t);\ldots;\nu_{n}(t)]$. Now by considering individual subsystems as $\Phi_{i}$ as
$\Phi_{i}\!\!:\!\left\{\begin{array}[]{rl}\mathsf{d}x_{i}(t)\!=\!\!&(a_{1}x_{i}^{3}(t)\!+\!b_{1}\nu_{i}(t)\!+\!d_{1}w_{i}(t))\mathsf{d}t+0.6\mathsf{d}\mathbb{W}_{t}+~{}0.5\mathsf{d}\mathbb{P}_{t},\quad\quad\quad t\in\mathbb{R}_{0}^{+}\backslash\Delta_{i},\\ x_{i}(t)\!=\!&a_{2}x_{i}^{3}(t^{-})\!+\!b_{2}\nu_{i}(t)\!+\!d_{2}w_{i}(t)\!+\!0.5\varsigma_{i}(t),\quad\quad\quad\quad\quad\quad\quad\quad\quad\!\!\!\!~{}t\!\in\!\Delta_{i},\end{array}\right.$ (7.6)
with $a_{1},b_{1},a_{2},b_{2}\in\mathbb{R}$, jump parameters $\tau=0.1$, $\varepsilon_{1_{i}}=1,\varepsilon_{2_{i}}=7$, and rates of Poisson processes as $\lambda_{i}=0.5$, one can readily verify that $\Phi=\mathcal{I}(\Phi_{1},\ldots,\Phi_{N})$ where $w_{i}(t)=x_{i\pm 1}(t)$ (with $x_{0}=x_{N}$ and $x_{N+1}=x_{1}$), i.e., circular interconnection topology.
The regions of interest are given as $X\in[0,8]^{1000},X_{0}\in[0,1.5]^{1000},X_{u}=[7,8]^{1000}$. The main goal is to design an A-CBC for the interconnected A-SHS and its corresponding safety controllers based on A-CSBCs of individual subsystems such that the state of the system remains in the comfort zone $[0,7]^{1000}$. To do so, we first search for CBCs and accordingly design local controllers for $\Phi_{i}$. Consequently, A-CSBCs can be constructed from CBCs according to Theorem 5.3. We show our results for different values of $a_{1},b_{1},a_{2},b_{2}$.
We employ the toolbox SOSTOOLS [PAV+13] and the SDP solver SeDuMi [Stu99] to compute CBC as described in Section 6. For $a_{1}=-0.4,b_{1}=0.5,a_{2}=0.01,b_{2}=0.06$, based on Lemma 6.1, we compute a CBC of an order $4$ as $\bar{\mathcal{B}_{i}}(x_{i})=0.0049x_{i}^{4}-0.0311x_{i}^{3}+0.0721x_{i}^{2}-0.0734x_{i}+0.0557$ and corresponding controllers $\nu_{i}=-0.05152x_{i}+3$ and $\hat{\nu}_{i}=-0.06145x_{i}+2.6$ for flow and jump cases, respectively. Moreover, the corresponding constants in Definition 5.1 satisfying conditions (5.1a)-(5.1e) are quantified as $\bar{\mu}_{i}=0.13,~{}\bar{\beta}_{i}=4.4,~{}\bar{\kappa}_{i}=0.005$, $\bar{\eta}_{1_{i}}=0.0015,~{}\bar{\eta}_{2_{i}}=0.0012,~{}\bar{\varphi}_{1_{i}}=0.00001,~{}\bar{\varphi}_{2_{i}}=0.00001,~{}\gamma_{1_{i}}=0.01,~{}\gamma_{2_{i}}=0.99$. We now proceed with Theorem 5.3 to construct an A-CSBC for the augmented system using the obtained CBC. We select $\alpha_{1_{i}}=0.1,\alpha_{2_{i}}=8$. Since $\gamma_{1_{i}}=0.01>0$ and $10<\gamma_{2_{i}}=0.99<$ 1, the results of Bound 1 are useful. One can readily verify that condition (5.3) is also met for all $\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\}$. Then one can conclude that $\mathcal{B}_{i}(x_{i},\vartheta_{i}):=\bar{\mathcal{B}}_{i}(x_{i})$ is a CSBC for the augmented SHS with $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau},\gamma_{2_{i}}\}=0.99$, $\varphi_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}=0.00001$ and $\eta_{i}\!=\!\max\{e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}=0.0012$. We now proceed to construct an A-CBC based on A-CSBCs of individual subsystems according to compositionality Theorem 4.2. By selecting $\xi_{i}=0.001$$\forall i\in\{1,\dots,1000\}$, one has $\xi_{i}\Gamma_{i}=-6\times\!10^{-6}\!<\!0,$ implying the satisfaction of condition (4.1a). Moreover, we compute $\beta=\sum_{i=1}^{1000}\xi_{i}\beta_{i}=4.4$, and $\mu=\sum_{i=1}^{1000}\xi_{i}\mu_{i}=0.13,$ that imply condition (4.1b). Then, it can be concluded that, $\mathcal{B}(x,\vartheta)=\sum_{i=1}^{1000}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i}),$ satisfying conditions in (3.2) with $\eta=\sum_{i=1}^{N}\xi_{i}\eta_{i}=0.0012,$ and $\gamma=1+\Gamma=0.994$.
Figure 3. Control barrier certificate of a representative subsystem within the interconnected SHS $\Phi$ for $a_{1}=-0.4,b_{1}=0.5,a_{2}=0.01,b_{2}=0.06$ (left), $a_{1}=-0.3,b_{1}=0.2,a_{2}=1.01,b_{2}=1$ (middle), $a_{1}=0.01,b_{1}=0.7,a_{2}=0.02,b_{2}=0.9$ (right). Initial and unsafe regions are depicted by green  and red  boxes. Moreover, ¯$\bar{\mathcal{B}}_{i}(x_{i})=\bar{\mu}_{i}$ and ¯$\bar{\mathcal{B}}_{i}(x_{i})=\bar{\beta}_{i}$ indicated by  and  , respectively.
For $a_{1}=-0.3,b_{1}=0.2,a_{2}=1.01,b_{2}=1$, we compute a CBC of an order $4$ as ¯$\bar{\mathcal{B}}_{i}(x_{i})=0.006295x^{4}_{i}-0.04521x^{3}_{i}+0.1213x^{2}_{i}-0.1449x_{i}+0.06662$ and corresponding controllers $\nu_{i}=-0.02152x^{3}_{i}+4$ (flow case) and $\hat{\nu}_{i}=-0.99x^{3}_{i}+2$ (jump case). Furthermore, the corresponding constants satisfying conditions (5.1a)-(5.1e) are obtained as $\bar{\mu}_{i}=0.12,~{}\bar{\beta}_{i}=4.6,~{}\bar{\kappa}_{i}=0.0043,$ $\gamma_{1_{i}}=0.04547,\gamma_{2_{i}}=1.00001,~{}\bar{\varphi}_{1_{i}}=1.7881\times 10^{-5},\bar{\varphi}_{2_{i}}=1.4408\times 10^{-5}$, $\bar{\eta}_{1_{i}}=0.0028,~{}\bar{\eta}_{2_{i}}=0.003$. Since $\gamma_{1_{i}}=0.04547>0$ and $\gamma_{2_{i}}=1.00001\geq 1$, Bound 2 is valid. Condition (5.3) is also met for all $\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\}$. Then $\mathcal{B}_{i}(x_{i},\vartheta_{i}):=e^{4.54\vartheta_{i}}\bar{\mathcal{B}}_{i}(x_{i})$ is a CSBC for the augmented SHS with $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})},e^{-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}}\gamma_{2_{i}}\}=0.9996,\varphi_{i}=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}=1.4408\times 10^{-5}$, and $\eta_{i}=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}$ $e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}=0.003$. We now proceed to construct an A-CBC based on A-CSBCs of individual subsystems according to compositionality Theorem 4.2. By choosing $\xi_{i}=0.001,$ $\forall i\in\{1,\dots,1000\}$, one has $\xi_{i}\Gamma_{i}=-2.9745\times\!10^{-6}\!<\!0,$ implying the satisfaction of condition (4.1a). Moreover, we compute $\beta=\sum_{i=1}^{1000}\xi_{i}\beta_{i}=4.6$, and $\mu=\sum_{i=1}^{1000}\xi_{i}\mu_{i}=0.12,$ that imply the satisfaction of condition (4.1b). Hence, it can be concluded that $\mathcal{B}(x,\vartheta)=\sum_{i=1}^{1000}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})$, implying the satisfaction of conditions in (3.2) with $\eta=\sum_{i=1}^{N}\xi_{i}\eta_{i}=0.003,$ and $\gamma=1+\Gamma=0.997.$
For $a_{1}=0.01,b_{1}=0.7,a_{2}=0.02,b_{2}=0.9$, we compute a CBC of an order $4$ as ¯$\bar{\mathcal{B}}_{i}(x_{i})=4.868\times 10^{-11}x^{6}_{i}+2.007\times 10^{-7}x^{5}_{i}+0.006941x^{4}_{i}-0.05582x^{3}_{i}+0.1685x^{2}_{i}-0.2275x_{i}+0.1563$ and its corresponding controllers $\nu_{i}=-0.2852x^{3}_{i}+2.5$ (flow case) and $\hat{\nu}_{i}=-0.09x^{3}_{i}+3$ (jump case). In addition, the corresponding constants in Definition 5.1 are acquired as $\bar{\mu}_{i}=0.16,~{}\bar{\beta}_{i}=4.2,~{}\bar{\kappa}_{i}=0.0043,\gamma_{1_{i}}=-0.0005,~{}\gamma_{2_{i}}=0.98$, $\bar{\varphi}_{1_{i}}=1.7881\times 10^{-5},~{}\bar{\varphi}_{2_{i}}=1.4408\times 10^{-5}$, and $\bar{\eta}_{1_{i}}=0.004,~{}\bar{\eta}_{2_{i}}=0.004$. Since $\gamma_{1_{i}}=-0.0005\leq 0$ and $10<\gamma_{2_{i}}=0.98<$ 1, Bound 3 is useful. Condition (5.3) is also met for all $\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\}$. Consequently, one can conclude that $\mathcal{B}_{i}(x_{i},\vartheta_{i}):=0.98^{\frac{\vartheta_{i}}{8}}\bar{\mathcal{B}_{i}}(x_{i})$ is a CSBC for the augmented SHS with $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}},\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}\}=0.9975$, $\varphi_{i}=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}=1.4408\times 10^{-5}$, and $\eta_{i}=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}=0.004$. We now proceed to construct an A-CBC based on A-CSBCs of individual subsystems according to compositionality Theorem 4.2. By selecting $\xi_{i}=0.001$, $\forall i\in\{1,\dots,1000\}$, one has $\xi_{i}\Gamma_{i}=-3.3007\times\!10^{-6}\!<\!0,$ implying the satisfaction of condition (4.1a). Moreover, we compute $\beta=\sum_{i=1}^{1000}\xi_{i}\beta_{i}=4.2$, and $\mu=\sum_{i=1}^{1000}\xi_{i}\mu_{i}=0.16,$ that imply the fulfillment of condition (4.1b). Then, it can be concluded that $\mathcal{B}(x,\vartheta)=\sum_{i=1}^{1000}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})$, satisfying conditions in (3.2) with $\eta=\sum_{i=1}^{N}\xi_{i}\eta_{i}=0.004,$ and $\gamma=1+\Gamma=0.9967.$
Figure 4. Closed-loop state trajectories of a representative subsystem within the interconnected SHS $\Phi$ with $10$ noise realizations for $a_{1}=-0.4,b_{1}=0.5,a_{2}=0.01,b_{2}=0.06$ (left), $a_{1}=-0.3,b_{1}=0.2,a_{2}=1.01,b_{2}=1$ (middle), $a_{1}=0.01,b_{1}=0.7,a_{2}=0.02,b_{2}=0.9$ (right).
By employing Theorem 3.5, we guarantee that the state trajectory of $\Phi$ starting from initial conditions inside $X_{0}=[0~{}1.5]^{1000}$ remains in the safe set $[0~{}7]^{1000}$ during the time steps $\mathcal{T}=100$ with the probability of at least $0.973$ (first bound), $0.9133$ (second bound), and $0.873$ (last bound). A control barrier certificate for a representative subsystem within the interconnected SHS $\Phi$ is depicted in Fig. 3. Closed-loop state trajectories of the interconnected SHS $\Phi$ for different bounds with $10$ noise realizations are depicted in Fig. 4. Note that we have only plotted trajectories at the discrete time steps since our guarantee is provided for them. The computation of CBC and its corresponding controllers for each bound took almost $5.3$ seconds with a memory usage of $64.94~{}\mathsf{Mbit}$ on a MacBook Pro M2 Max chip with 32 GB of memory.
8. Conclusion
In this work, we proposed a compositional framework for the safety controller synthesis of interconnected stochastic hybrid systems with both continuous and discrete evolutions. In our proposed setting, we first introduced an augmented framework to reformulate each stochastic hybrid subsystem with continuous and discrete evolutions with a unified framework covering both behaviours. We then introduced notions of augmented control sub-barrier certificates, constructed for each subsystem, by employing which we constructed augmented control barrier certificates for interconnected systems together with their safety controllers under some small-gain compositional conditions. By employing the constructed augmented control barrier certificate, we finally computed a guaranteed probabilistic bound on the safety of the interconnected stochastic hybrid systems. We verified the effectiveness of our proposed approaches by applying them to an interconnected system composed of $1000$ nonlinear subsystems. Developing a compositional approach for stochastic hybrid systems capable of handling more complex logical properties beyond safety is being investigated as future work.