### v2

# Compositional Design of Safety Controllers for Large-scale Stochastic Hybrid Systems

Abstract
In this work, we propose a compositional scheme based on small-gain reasoning for the safety controller synthesis of interconnected stochastic hybrid systems with both continuous evolutions and instantaneous jumps. In our proposed setting, we first offer an augmented scheme to represent each stochastic hybrid subsystem with continuous and discrete evolutions in a unified framework, ensuring that the state trajectories match those of the original hybrid systems. We then introduce the concept of augmented control sub-barrier certificates (A-CSBC) for each subsystem, which allows the construction of augmented control barrier certificates (A-CBC) for interconnected systems and their safety controllers under small-gain compositional conditions. We eventually leverage the constructed A-CBC and quantify a guaranteed probabilistic bound across the safety of the interconnected system. While the computational complexity of designing a barrier certificate and its safety controller grows polynomially with network dimension using sum-of-squares (SOS) optimization program, our compositional approach significantly reduces it to a linear scale with respect to the number of subsystems. We verify the efficacy of our proposed approach over an interconnected stochastic hybrid system composed of $1000$ nonlinear subsystems with two distinct interconnection topologies.
keywords:
Stochastic hybrid systems, scalable safety controllers, control barrier certificates, small-gain reasoning
, , , and
1 Introduction
Stochastic hybrid systems (SHS) (Blom et al., 2006; Cassandras & Lygeros, 2006) are complex heterogeneous models, spanning a wide range of safety-critical applications, in which discrete dynamics describe computation and communication parts and continuous models capture physical components. Providing a formal verification and controller synthesis framework for such complex heterogeneous systems is naturally very challenging primarily due to (i) the stochastic nature of dynamics, (ii) the tight interaction between continuous and discrete components (in both space and time), (iii) the large dimension of state/input sets, and (iv) the need to manage complex logic requirements (Baier & Katoen, 2008).
Discretization-based approaches. In order to deal with the above-mentioned difficulties, one potential solution, proposed in the relevant literature, is to approximate original models by a simpler one with either a lower dimension (a.k.a., infinite abstractions) or with discrete-state sets (a.k.a., finite abstractions) (Julius & Pappas, 2009; Abate et al., 2008). On the downside, the existing abstraction-based techniques rely on discretizing state/input sets and are not applicable to real-world high-dimensional systems due to the state-explosion problem. To make the abstraction-based techniques more scalable, compositional approaches have been proposed in the relevant literature that construct abstractions of large-scale systems based on those of smaller subsystems (Lavaei & Zamani, 2022; Hahn et al., 2013; Nejati et al., 2020; Lavaei, Soudjani, Abate & Zamani, 2022; Lavaei, 2019).
Discretization-free approaches. To mitigate the computational complexity arising from abstraction-based approaches, control barrier certificates (CBC), a promising discretization-free approach, have garnered significant attention in recent years for the formal analysis of complex dynamical systems. Specifically, barrier certificates are Lyapunov-like functions whose initial level set separates an unsafe region from all system trajectories starting from a set of initial states. Consequently, the existence of such a function provides a (probabilistic) safety certificate over the system. This approach has been initially proposed by Prajna & Jadbabaie (2004); Prajna et al. (2007) and then widely employed for formal verification and synthesis of non-stochastic (Borrmann et al., 2015; Ames et al., 2019; Wooding et al., 2025) and stochastic dynamical systems (Zhang et al., 2010; Jahanshahi et al., 2022; Ahmadi et al., 2019; Nejati et al., 2024; Santoyo et al., 2019; Clark, 2019; Nejati & Zamani, 2022; Lavaei & Frazzoli, 2024; Nejati & Schmuck, 2024), to name a few.
Innovative findings. Despite the extensive above-mentioned literature on barrier certificate construction, there are no results available for large-scale stochastic hybrid systems. Our primary contribution is to propose a compositional scheme for the safety controller synthesis of interconnected SHS with both continuous evolutions, governed by stochastic differential equations with Brownian motions and Poisson processes, and instantaneous jumps, modeled by discrete-time stochastic reset equations with additive noises. In our proposed setting, we first introduce an augmented framework to recast each stochastic hybrid subsystem with continuous and discrete evolutions via a unified framework, spanning both scenarios. We then introduce notions of augmented control sub-barrier certificates, computed for each subsystem, using which one can construct augmented control barrier certificates for interconnected systems together with their safety controllers under some small-gain compositional conditions. We eventually leverage the constructed augmented control barrier certificate and offer a certified probabilistic bound across the safety of the interconnected system.
While a recent work by Lavaei & Frazzoli (2024) addresses networks of discrete-time switched systems with jumps acting solely through the transition map, our work considers a broader class of stochastic hybrid systems. These include both continuous- and discrete-time (reset) dynamics, with switching in both flow and jump components, modeled via stochastic differential and reset equations driven by Brownian motion, Poisson processes, and additive noise. This richer structure requires safety analysis for both evolution modes and necessitates additional design elements (see Theorem 5.4) to ensure soundness across the hybrid dynamics. Moreover, unlike (Lavaei & Frazzoli, 2024), which lacks control inputs and focuses on switching signal design under dwell-time constraints, our model explicitly incorporates control inputs in both the flow and jump dynamics. Accordingly, the controller synthesis problem becomes more challenging in our setting, as it involves two distinct types of adversarial inputs: (i) disturbance inputs, which capture the influence of other subsystems, and (ii) switching between flow and jump modes in a hybrid setting with both continuous- and discrete-time dynamics. Finally, while Lavaei & Frazzoli (2024) adopts a dissipativity-based compositional method, our approach is grounded in small-gain reasoning, which demands carefully designing the disturbance gains to be sufficiently small to enable lifting safety guarantees from individual subsystems to the entire interconnected network. This is particularly challenging in stochastic hybrid systems due to their hybrid nature.
While the primary results of this work were initially presented in (Lavaei, Soudjani & Frazzoli, 2022), they are limited to monolithic systems and cannot handle models with more than six state dimensions due to the computational complexity involved in searching for control barrier certificates. In contrast, we propose here an innovative compositional technique to address large-scale SHS by constructing an augmented control barrier certificate for the interconnected system based on certificates of individual subsystems (cf. the $1000$-dimensional case study). While the computational complexity of designing a barrier certificate and its safety controller in (Lavaei, Soudjani & Frazzoli, 2022) grows polynomially with system dimension using sum-of-squares (SOS) optimization program, our compositional approach here significantly reduces it to a linear scale with respect to the number of subsystems (cf. Subsection 7.4). Moreover, a key challenge in analyzing networked stochastic hybrid systems lies not only in the high dimensionality inherent to such systems, but also in the coupling between subsystems. This coupling, modeled as the adversarial input in each subsystem’s dynamics (see (2.4)), poses a central difficulty specifically due to the hybrid nature of the system, necessitating a more careful treatment of these interconnections (cf. Remark 5.6), which is absent in (Lavaei, Soudjani & Frazzoli, 2022). Additionally, we present the proofs for all statements that were not included in (Lavaei, Soudjani & Frazzoli, 2022).
2 Stochastic Hybrid Systems
2.1 Notation
Sets of non-negative and positive integers are, respectively, denoted by $\mathbb{N}:=\{0,1,2,$...\} and $\mathbb{N}^{+}=\{1,2,$...\}. Sets of real, positive and non-negative real numbers are represented by $\mathbb{R},\mathbb{R}^{+}$, and $\mathbb{R}_{0}^{+}$, respectively. Given $N$ vectors $_{i}\in\mathbb{R}^{n_{i}}, x=[x$$x_{1};...;x_{N}$] represents the corresponding vector of dimension $\sum_{i}n_{i}$. Given a vector $x\in\mathbb{R}^{n}$ , we define the Hadamard power as $x^{\circ m}=[x_{1}^{m}~...~x_{n}^{m}]^{\top}$ for $m\in\mathbb{N}$. A (block) diagonal matrix in ${\mathbb{R}}^{N\times{N}}$ with diagonal matrix entries $(A_{1},\ldots,A_{N})$ and scalar entries $(a_{1},\ldots,a_{N})$ is denoted by $\mathsf{blkdiag}(A_{1},\ldots,A_{N})$ and $\mathsf{diag}(a_{1},\ldots,a_{N})$, respectively. The notation $\{a_{ij}\}$ denotes a matrix whose entry in the $i$-th row and $j$-th column is given by $a_{ij}$. We denote the Euclidean norm of a vector $x\in\mathbb{R}^{n}$ by $\|x\|$. Given sets $X_{i},i\in\{1,\ldots,N\}$, their Cartesian product is denoted by $\prod_{i=1}^{N}X_{i}$. The empty set is represented by $\emptyset$. Symbol $\mathds{I}_{n}$ denotes the identity matrix in $\mathbb{R}^{n\times{n}}$. Given a measurable function $f:\mathbb{N}\rightarrow\mathbb{R}^{n}$, the (essential) supremum of $f$ is denoted by $\|f\|_{\infty}:=\text{(ess)sup}\{\|f(k)\|,k\geq 0\}$.
2.2 Preliminaries
Given a probability space $(\Omega,\mathcal{F}_{\Omega},\mathds{P}_{\Omega})$, with $\Omega$ being the sample space, $\mathcal{F}_{\Omega}$ a sigma-algebra on $\Omega$, and $\mathds{P}_{\Omega}$ a probability measure, we assume that $(\Omega,\mathcal{F}_{\Omega},\mathds{P}_{\Omega})$ is endowed with a filtration $\mathds{F}=(\mathcal{F}_{s})_{s\geq 0}$ fulfilling conditions of completeness and right continuity. We consider $(\mathbb{W}_{s})_{s\geq 0}$ as a b-dimensional $\mathds{F}$-Brownian motion and $(\mathbb{P}_{s})_{s\geq 0}$ as an r-dimensional $\mathds{F}$-Poisson process. The Poisson process and Brownian motion are considered to be mutually independent. The Poisson process $\mathbb{P}_{s}=[\mathbb{P}_{s}^{1};\cdots;\mathbb{P}_{s}^{\textsf{r}}]$ models r events with independent occurrences.
2.3 Stochastic Hybrid Systems
In this work, our focus lies on stochastic hybrid systems, treating them as individual subsystems. We begin by formally introducing the notion of a stochastic hybrid system (SHS), which generalizes the impulsive system concept in (Swikir et al., 2020, Definition 1) by including different sources of noises and disturbances.
Definition 2.1.
A stochastic hybrid system (SHS) $\Phi_{i}$ is characterized by $\Phi_{i}\!=\!(X_{i},U_{i},\mathcal{U}_{i},W_{i},\mathcal{W}_{i},\sigma_{i},\rho_{i},f_{1_{i}},$ $\varsigma_{i},f_{2_{i}})$, where:
  • $X_{i}\subseteq\mathbb{R}^{n_{i}}$ is the state set of SHS;
  • $U_{i}\subseteq\mathbb{R}^{m_{i}}$ is the input set of SHS;
  • $\mathcal{U}_{i}$ is the set of all measurable bounded input functions $\nu_{i}:\mathbb{R}_{0}^{+}\rightarrow U_{i}$;
  • $W_{i}\subseteq\mathbb{R}^{p_{i}}$ is the disturbance set of SHS (which is utilized later for the sake of interconnection);
  • $\mathcal{W}_{i}$ is the set of all measurable bounded disturbance functions $w_{i}:\mathbb{R}_{0}^{+}\rightarrow W_{i}$;
  • $\sigma_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{i}\times\textsf{b}_{i}}$ is the diffusion term which is globally Lipschitz continuous;
  • $\rho_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{i}\times\textsf{r}_{i}}$ is the reset term which is globally Lipschitz continuous;
  • $f_{1_{i}}:X_{i}\times U_{i}\times W_{i}\rightarrow X_{i}$ is the drift term which is globally Lipschitz continuous;
  • $\varsigma_{i}$ is a sequence of independent and identically distributed (i.i.d.) random variables from the sample space $\Omega_{i}$ to the measurable space $(\mathcal{V}_{\varsigma_{i}},\mathcal{F}_{\varsigma_{i}})$:
    $\varsigma_{i}:=\Big\{\varsigma_{i}(\cdot)\!:(\Omega_{i},\mathcal{F}_{\Omega_{i}})\rightarrow(\mathcal{V}_{\varsigma_{i}},\mathcal{F}_{\varsigma_{i}})\Big\};$
  • $f_{2_{i}}:X_{i}\times U_{i}\times W_{i}\times\mathcal{V}_{\varsigma_{i}}\rightarrow X_{i}$ is the transition map which is globally Lipschitz continuous.
The SHS $\Phi_{i}$ can be characterized by the following stochastic differential and reset equations:
$\Phi_{i}\!:\left\{\hskip-4.2679pt\begin{array}[]{rl}\mathsf{d}x_{i}(t)\!=\!\!&f_{1_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t))\mathsf{d}t\!+\!\sigma_{i}(x_{i}(t))\mathsf{d}\mathbb{W}_{t}\\ &+\rho_{i}(x_{i}(t))\mathsf{d}\mathbb{P}_{t},~\quad\quad\quad\quad\quad\quad~~~t\in\mathbb{R}_{0}^{+}\backslash\Delta,\\ x_{i}(t)\!=\!\!&{f_{2_{i}}(x_{i}(t^{-}),\nu_{i}(t),w_{i}(t^{-}),\varsigma_{i}(t))},\quad\!\!\!t\in\Delta,\end{array}\right$. (2.4)
where $\Delta=\{t_{k}\}_{k\in\mathbb{N}}$ with $t_{k+1}-t_{k}\in\{\varepsilon_{1_{i}}\tau,\ldots,\varepsilon_{2_{i}}\tau\}$ for jump parameters $\tau\in\mathbb{R}^{+}$ and $\varepsilon_{1_{i}},\varepsilon_{2_{i}}\in\mathbb{N}^{+}$, $\varepsilon_{1_{i}}\leq\varepsilon_{2_{i}}$. Here, $\mathbb{W}_{t}$ denotes the Brownian motion, and $\mathbb{P}_{t}$ represents the Poisson process with rate $\lambda$. Moreover, $f_{2_{i}}(x_{i}(t^{-}),\nu_{i}(t),w_{i}(t^{-}),\varsigma_{i}(t))=\lim_{t\to t_{k}^{-}}f_{2_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t),\varsigma_{i}(t))$, indicating the value of $f_{2_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t),\varsigma_{i}(t))$ as $t$ goes to $t_{k}$ from left. Furthermore, $x_{i}:\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}^{n_{i}}$ is the state signal (right-continuous for all $t\in\mathbb{R}_{0}^{+}$) and $\nu_{i}(\cdot)\in\mathcal{U}_{i},w_{i}(\cdot)\in\mathcal{W}_{i}$ are input and disturbance signals. The random sequence $x_{x_{0_{i}}\nu_{i}w_{i}}(t)$ fulfilling (2.4) is called the state trajectory (a.k.a. solution) of $\Phi_{i}$ at time $t\in\mathbb{R}_{0}^{+}$ under signals $\nu_{i}(\cdot),w_{i}(\cdot)$ starting from $x_{0_{i}}$.
Remark 2.2.
In the SHS defined in Definition 2.1, jumps at time instants $t\in\Delta$ introduce discontinuities in the state $x_{i}(t)$, which evolves discretely at such times. To ensure the well-posedness of the model, it is necessary to specify the left-hand limit $x_{i}(t^{-})$ at each jump instant. The same requirement applies to the adversarial inputs $w_{i}(t)$, which represent the (possibly discontinuous) states of neighboring subsystems (cf. (2.7)). In contrast, the control input $\nu_{i}(t)$ is user-designed and can be applied as a continuous-time signal without needing to exhibit jumps (Swikir et al., 2020). The additive noise $\varsigma_{i}(t)$ is also treated as externally imposed and does not require specification at $t^{-}$ as it is not necessarily exhibiting jumps.
Remark 2.3.
The drift, diffusion, reset functions, and the transition map of the stochastic hybrid system are all assumed to be globally Lipschitz continuous. This assumption, stated in Definition 2.1, ensures the existence, uniqueness, and strong Markov property of the solution process $x_{x_{0_{i}}\nu_{i}w_{i}}(t)$, and thereby, the well-posedness of the stochastic hybrid system. Of note is, since our CBC analysis is restricted to compact domains, any continuous function defined on such domains is automatically globally Lipschitz continuous.
To further clarify the modeling capabilities of such systems and the roles of the Brownian motion and Poisson process, we present the following illustrative example.
Example 2.4.
Consider a network of $N$ rooms, each equipped with a local heater. The dynamics in (2.4) can model each room, where $x_{i}(t)$ denotes its temperature, controlled by the heater control input $\nu_{i}(t)$, and influenced by $w_{i}(t)$, the inter-room heat transfer (e.g., through shared walls). The Brownian motion $\mathbb{W}_{t}$ models continuous environmental uncertainties (e.g., brief door/window openings), while the Poisson process $\mathbb{P}_{t}$ captures discrete events like people entering, causing sudden temperature changes. The jump times $\Delta=\{t_{k}\}_{k\in\mathbb{N}}$ correspond to sudden thermostat reconfigurations (e.g., switching between ECO and COMFORT modes), occurring at intervals constrained by $\varepsilon_{1_{i}}\tau\leq t_{k+1}-t_{k}\leq\varepsilon_{2_{i}}\tau$.
Given that control inputs in practice are often fed by a digital controller, we are interested here in sampled-data hybrid systems in which the control inputs are considered to be piecewise constant as
$\mathcal{U}_{\tau_{i}}\!\!=\!\Big\{\nu_{i}\!:\mathbb{R}_{0}^{+}\!\rightarrow\!U_{i}\,\big|\,\nu_{i}(t)\!=\!\nu_{i}(t_{k-1}),{t\!\in\![t_{k-1},t_{k}),k\!\in\!\mathbb{N}^{+}}\!\!\Big\},$ (2.5)
with $t_{k}=k\tau$.
Remark 2.5.
While our theoretical framework does not require the control inputs $\nu_{i}$ to be piecewise constant, we adopt this assumption as control inputs in digital practical implementations are typically applied using a zero-order hold mechanism, which naturally results in piecewise constant signals.
2.4 Interconnected SHS
Since the ultimate goal in this work is to synthesize a safety controller for an interconnected SHS, we provide next a formal definition for interconnected SHS without disturbance signals $w$ that can be constructed as a composition of multiple SHS with disturbance signals.
Definition 2.6.
Consider $N\!\in\!\mathbb{N}^{+}$ SHSs $\Phi_{i}=(X_{i},U_{i},$ $\mathcal{U}_{i},W_{i},\mathcal{W}_{i},\sigma_{i},\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$, with their disturbances partitioned as
$w_{i}$ $=[{w_{i1};\ldots;w_{i(i-1)};w_{i(i+1)};\ldots;w_{iN}}].$ (2.6)
The interconnection of $\Phi_{i}$ is $\Phi=(X,U,\mathcal{U},\sigma,\rho,f_{1},\varsigma,f_{2})$, denoted by $\mathcal{I}(\Phi_{1},\ldots,\Phi_{N})$, such that $X=\Pi_{i=1}^{N}X_{i},U=\Pi_{i=1}^{N}U_{i},\sigma:=\mathsf{blkdiag}(\sigma_{1}(x_{1}),\ldots,\sigma_{N}(x_{N})),\rho:=\mathsf{blkdiag}\\ (\rho_{1}(x_{1}),\ldots,\rho_{N}(x_{N})),f_{1}:=[f_{1_{1}};\ldots;f_{1_{N}}],f_{2}:=[f_{2_{1}};\ldots;f_{2_{N}}]$ and $\varsigma:=[\varsigma_{1};\dots;\varsigma_{N}]$, subjected to the following interconnection constraint:
$\forall i,j\in\{1,\dots,N\},i\neq j\!:~~w_{ij}=x_{j},~~X_{j}\subseteq W_{ij},$ (2.7)
where $W_{i}:=\prod_{j}W_{ij}$. Such an interconnected SHS can be described by
$\Phi\!:\left\{\hskip-4.2679pt\begin{array}[]{rl}\mathsf{d}x(t)=\!\!&f_{1}(x(t),\nu(t))\mathsf{d}t+\sigma(x(t))\mathsf{d}\mathbb{W}_{t}+\rho(x(t))\mathsf{d}\mathbb{P}_{t},\\ &~\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad~t\in\mathbb{R}_{0}^{+}\backslash\Delta,\\ x(t)=\!\!&{f_{2}(x(t^{-}),\nu(t),\varsigma(t))},\,\quad\quad\quad\quad\quad~\!\!t\in\Delta.\end{array}\right$. (2.11)
The random sequence $x_{x_{0}\nu}(t)$ fulfilling (2.11) is called the state trajectory of $\Phi$ at time $t\in\mathbb{R}_{0}^{+}$ under an input signal $\nu(\cdot)$ starting from $x_{0}$.
We now formally define the safety specification of interest for the interconnected SHS.
Definition 2.7.
Given a safety specification $\Lambda=(X_{0},X_{u},\mathcal{T})$, with $X_{0},X_{u}\subseteq X$ representing the initial and unsafe sets of the SHS $\Phi$, where $X_{0}\cap X_{u}=\emptyset$, the system $\Phi$ is said to be safe within the time horizon $\mathcal{T}\in\mathbb{N}$, denoted by $\Phi\vDash\Lambda$, if all trajectories starting from $X_{0}$ avoid $X_{u}$ during $\mathcal{T}$. Since the trajectories of $\Phi$ are stochastic, we are interested in computing the probability $\mathds{P}\{\Phi\vDash\Lambda\}\geq 1-\epsilon$, for some $\epsilon\in(0,1]$.
2.5 Augmented SHS
As can be observed, the SHS in (2.4) encompasses both continuous and discrete evolutions, in which continuous behaviors are governed by stochastic differential equations with Brownian motions and Poisson processes, and instantaneous jumps are modeled by discrete-time stochastic reset equations with additive noises. To enable a unified treatment of these two modes of evolution—and to handle the logic governing jump timing—we introduce an augmented representation of SHS, adopted from (Swikir et al., 2020, Def. 5) but adapted to the stochastic setting, as formalized in the following definition.
Definition 2.8.
Consider an SHS $\Phi_{i}=(X_{i},U_{i},\mathcal{U}_{i},W_{i},\\ \mathcal{W}_{i},\sigma_{i},\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$ with jump parameters ($\tau$, $\varepsilon_{1_{i}}$, $\varepsilon_{2_{i}}$). An augmented SHS (A-SHS) $\mathcal{A}_{i}(\Phi_{i})$ associated to $\Phi_{i}$ is characterized by $\mathcal{A}_{i}(\Phi_{i})\!=\!(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},$ $\mathbb{H}_{i})$, where:
  • $\mathbb{X}_{i}=X_{i}\times\{0,\ldots,\varepsilon_{2_{i}}\}$ is the state space of A-SHS, where $(x_{i},\vartheta_{i})\in\mathbb{X}_{i}$ indicates that the current state is $x_{i}$, and the time elapsed since the latest jump is $\vartheta_{i}$;
  • $\mathbb{U}_{i}=\mathcal{U}_{\tau_{i}}$ is the input space of A-SHS;
  • $\mathds{W}_{i}$ is the disturbance space of A-SHS;
  • $\sigma_{i}$ is the diffusion term;
  • $\rho_{i}$ is the reset term;
  • $\varsigma_{i}$ is a sequence of i.i.d. random variables;
  • $(x_{i}^{\prime},\vartheta_{i}^{\prime})=\mathbb{F}_{i}((x_{i},\vartheta_{i}),\nu_{i},w_{i},\sigma_{i},\rho_{i},\varsigma_{i})$ if and only if one of the following two cases holds:
    1. (i)
      Flow case: $0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1$,
      $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, and $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
    2. (ii)
      Jump case: $\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}}$,
      $x_{i}^{\prime}=f_{2_{i}}(x_{i}(\tau^{-}),\nu_{i}(\tau),w_{i}(\tau^{-}),\varsigma_{i}(\tau))$, and $\vartheta_{i}^{\prime}=0$;
  • $\mathbb{Y}_{i}=\mathbb{R}^{n_{i}}$ is the output space of A-SHS;
  • $\mathbb{H}_{i}:\mathbb{X}_{i}\rightarrow\mathbb{Y}_{i}$ is the output map defined as $\mathbb{H}_{i}(x_{i},\vartheta_{i})=x_{i}$.
Remark 2.9.
Note that $(x_{i}^{\prime},\vartheta_{i}^{\prime})=\mathbb{F}_{i}((x_{i},\vartheta_{i}),\nu_{i},w_{i},\sigma_{i},\\ \rho_{i},\varsigma_{i})$ in Definition 2.8 characterizes the evolution of A-SHS, in which $\mathbb{F}_{i}$ is acquired based on one of the following two cases:
  • If the counter $\vartheta_{i}$ is between 0 and $\varepsilon_{2_{i}}-1$, then the state may evolve based on the flow case as $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, in which $x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$ is the solution of the stochastic differential equation $\mathsf{d}x_{i}(t)=f_{1_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t))\mathsf{d}t\!+\!\sigma_{i}(x_{i}(t))\mathsf{d}\mathbb{W}_{t}+\rho_{i}(x_{i}(t))\mathsf{d}\mathbb{P}_{t},$ which is well defined in the literature (Øksendal & Sulem, 2005). In this case, the next counter is increased as $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
  • If the counter $\vartheta_{i}$ is between $\varepsilon_{1_{i}}$ and $\varepsilon_{2_{i}}$, then the state may evolve based on the jump case as $x_{i}^{\prime}={f_{2_{i}}(x_{i}(\tau^{-}),\nu_{i}(\tau),w_{i}(\tau^{-}),\varsigma_{i}(\tau))}$ and the counter is reset from zero as $\vartheta_{i}^{\prime}=0$.
We further clarify the evolution of the SHS and the A-SHS in Definitions 2.1 and 2.8, respectively, and how they are related to each other through the following illustrative example.
Example 2.10.
Consider the case where $\varepsilon_{1_{i}}=1$ and $\varepsilon_{2_{i}}=3$. Under Definition 2.1, the system evolves continuously according to the stochastic differential equation until a jump occurs. If the first jump happens at time $t_{k}$, then the next jump at $t_{k+1}$ must satisfy $t_{k+1}-t_{k}\in\{\tau,2\tau,3\tau\},$ meaning that successive jumps occur no sooner than $\tau$ and no later than $3\tau$ after the preceding jump. If no jump takes place at $\tau$ or $2\tau$ following $t_{k}$, a jump is guaranteed at $3\tau$. This pattern repeats for every pair of consecutive jumps.
In Definition 2.8, the auxiliary state variable $\vartheta_{i}$ acts as a counter functioning as memory to track the elapsed time between consecutive jumps, and it increments by one at each time step of length $\tau$. Hence, when $\varepsilon_{1_{i}}=1$ and $\varepsilon_{2_{i}}=3$, while $\vartheta_{i}$ (an integer value) remains within the inclusive range $\varepsilon_{1_{i}},\varepsilon_{2_{i}}]$, a jump may occur—specifically at $\vartheta_{i}=1$ (after $\tau$) or at $\vartheta_{i}=2$ (after $2\tau$). If no jump happens by the time $\vartheta_{i}$ reaches $\varepsilon_{2_{i}}=3$, a jump is then enforced at $3\tau$. Immediately following any jump, $\vartheta_{i}$ resets to zero, beginning a new count toward the next permissible jump interval. More precisely, the system begins in the flow case with $\vartheta_{i}=0$, after which the counter increments to one ($\vartheta_{i}=1$). Then, since $\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}}$, a jump might happen, and if so, the system evolves following the jump case. Otherwise, $\vartheta_{i}$ increases to two ($\vartheta_{i}=2$). If a jump does not happen again, $\vartheta_{i}$ becomes three ($\vartheta_{i}=3$). Since $0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1$ does not hold anymore, the system necessarily evolves according to the jump case, and $\vartheta_{i}$ resets to zero to count the time elapsed until the next jump. Thereby, Definitions 2.1 and 2.8 represent the same model.
Although the equivalency of Definitions 2.1 and 2.8 is intuitively demonstrated thorough Example 2.10, in the following proposition, we formally show that the state trajectories of an A-SHS and its original SHS are equivalent. Consequently, we leverage the A-SHS as a unified framework that encompasses both continuous evolutions and instantaneous jumps, providing a more manageable model.
Proposition 2.11.
The state trajectory of $\Phi_{i}$ defined in (2.4) can be uniquely mapped to a state trajectory of $\mathcal{A}_{i}(\Phi_{i})$ in Definition 2.8 and vice versa.
[…]
3 Augmented Control (Sub-)Barrier Certificates
In this section, we introduce notions of augmented control sub-barrier and barrier certificates for, respectively, augmented SHS with and without disturbance signals. We then utilize those notions and quantify a guaranteed probabilistic bound over the safety of the interconnected SHS.
The following definition formalizes the notion of an augmented control sub-barrier certificate (A-CSBC), which acts as a local certificate for each subsystem, borrowed from (Lavaei, Di Lillo, Censi & Frazzoli, 2022, Def. 3.3), but adapted to the SHS.
Definition 3.1.
Given an A-SHS $\mathcal{A}_{i}(\Phi_{i})=(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\\ \sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},\mathbb{H}_{i})$, consider $\mathbb{X}_{0_{i}}=X_{0_{i}}\times\{0\},~\mathbb{X}_{u_{i}}=X_{u_{i}}\times\{0,\ldots,\varepsilon_{2_{i}}\}$, as initial and unsafe regions of $\mathcal{A}_{i}(\Phi_{i})$, respectively, with $X_{0_{i}},X_{u_{i}}\subseteq X_{i}$ being initial and unsafe regions of $\Phi_{i}$. A function $\mathcal{B}_{i}:\mathbb{X}_{i}\to{\mathbb{R}}_{0}^{+}$ is said to be an augmented control sub-barrier certificate (A-CSBC) for $\mathcal{A}_{i}(\Phi_{i})$ if there exist $0<\gamma_{i}<1$, $\kappa_{i}\in{\mathbb{R}}^{+},\mu_{i},\beta_{i},\varphi_{i},\eta_{i}\in{\mathbb{R}}_{0}^{+}$, such that
$\mathcal{B}_{i}(x_{i},\vartheta_{i})\geq\kappa_{i}\|x_{i}\|^{2},\quad\quad\quad\quad~\!\forall(x_{i},\vartheta_{i})\in\mathbb{X}_{i},$ (3.1a)
$\mathcal{B}_{i}(x_{i},\vartheta_{i})\leq\mu_{i},\quad\quad\quad\quad\quad\quad\quad\!\!\!\!\forall(x_{i},\vartheta_{i})\in\mathbb{X}_{0_{i}},$ (3.1b)
$\mathcal{B}_{i}(x_{i},\vartheta_{i})\geq\beta_{i},\quad\quad\quad\quad\quad\quad\quad\!\!\!\forall(x_{i},\vartheta_{i})\in\mathbb{X}_{u_{i}},$ (3.1c)
and $\forall(x_{i},\vartheta_{i})\in\mathbb{X}_{i}$, $\exists\nu_{i}\in\mathbb{U}_{i}$, such that $\forall w_{i}=w_{i}(t_{k})\in\mathds{W}_{i}$, $(x_{i}^{\prime},\vartheta_{i}^{\prime})=\mathbb{F}_{i}((x_{i},\vartheta_{i}),\nu_{i},w_{i},\sigma_{i},\rho_{i},\varsigma_{i})$, and
$\mathds{E}\Big[\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\big|x_{i},\nu_{i},w_{i},\vartheta_{i}\Big]\!\leq\!\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\varphi_{i}\|w_{i}\|^{2}+\eta_{i},$ (3.1d)
where $\mathds{E}$ is the expected value w.r.t $\varsigma_{i}$ under the one-step transition of $\mathcal{A}_{i}(\Phi_{i})$.
In this definition, the key inequality (3.1d) expresses how the expected certificate value evolves under one-step transitions. The presence of $\varphi_{i}$ reflects the influence of interconnections, which becomes crucial in compositional safety analysis.
Remark 3.2.
The order of quantifiers used in condition (3.1d), specifically $\forall x_{i}\in X_{i}$, $\exists\nu_{i}\in U_{i}$, $\forall w_{i}\in W_{i}$, inherently facilitates the synthesis of fully-decentralized controllers for $\mathcal{A}_{i}(\Phi_{i})$. This is because the control input $\nu_{i}$ is independent of the internal input $w_{i}$, which represents the state information of other subsystems. As a result, this independence offers considerable flexibility in designing controllers for large-scale networks.
This remark highlights a key design advantage: the control input $\nu_{i}$ is independent of the internal coupling term $w_{i}$, thereby enabling fully decentralized controller synthesis. This structural property facilitates scalable control design without the need for global coordination among subsystems. We now provide a notion of augmented control barrier certificates for augmented SHS without disturbance signals, i.e., interconnected A-SHS (cf. Definition 4.1), adapted from (Lavaei, Di Lillo, Censi & Frazzoli, 2022, Def. 3.1).
Definition 3.3.
Given an interconnected A-SHS $\mathcal{A}(\Phi)=(\mathbb{X},\mathbb{U},\sigma,\rho,\varsigma,\mathbb{F},\mathbb{Y},\mathbb{H})$ with initial and unsafe regions $\mathbb{X}_{0},\mathbb{X}_{u}$, a function $\mathcal{B}:\mathbb{X}\to{\mathbb{R}}_{0}^{+}$ is said to be an augmented control barrier certificate (A-CBC) for $\mathcal{A}(\Phi)$ if there exist $0<\gamma<1$, $\mu,\beta,\eta\in{\mathbb{R}}_{0}^{+}$, with $\beta>\mu$, such that
$\mathcal{B}(x,\vartheta)\leq\mu,\quad\quad\quad\quad\quad\quad\quad\quad\!\!\forall(x,\vartheta)\in\mathbb{X}_{0},$ (3.2a)
$\mathcal{B}(x,\vartheta)\geq\beta,\quad\quad\quad\quad\quad\quad\quad\quad\!\!\forall(x,\vartheta)\in\mathbb{X}_{u},$ (3.2b)
and $\forall(x,\vartheta)\in\mathbb{X}$, $\exists\nu\in U$, such that $(x^{\prime},\vartheta^{\prime})=\mathbb{F}((x,\vartheta),\nu,\sigma,\rho,\varsigma)$, and
$\mathds{E}\Big[\mathcal{B}(x^{\prime},\vartheta^{\prime})\,\big|\,x,\nu,\vartheta\Big]\leq\gamma\mathcal{B}(x,\vartheta)+\eta.$ (3.2c)
The A-CBC defined in Definition 3.2 serves as a global safety certificate, where conditions (3.2a)–(3.2b) ensure separation between initial and unsafe sets, while (3.2c) ensures expected safety evolution under controlled dynamics.
Note that $\beta>\mu$ is required in Definition 3.3 in order to ensure there is no intersection between initial and unsafe sets of interconnected A-SHS. The next theorem, borrowed from (Kushner, 1967) but adapted to A-SHS, shows the usefulness of A-CBC by quantifying a guaranteed probabilistic bound on the safety of the interconnected SHS, as a function of the certificate bounds and decay parameters.
Theorem 3.4.
Given interconnected A-SHS $\mathcal{A}(\Phi)=(\mathbb{X},\mathbb{U},\sigma,\rho,\varsigma,\mathbb{F},\mathbb{Y},\mathbb{H})$, let $\mathcal{B}$ be an A-CBC for $\mathcal{A}(\Phi)$. Then for any $x_{0}$ and $\vartheta_{0}$, the probability that the state trajectories of $\mathcal{A}(\Phi)$ reach the unsafe region $\mathbb{X}_{u}$ within the finite-time horizon $k\in{[0,1,\dots\mathcal{T}]}$, is computed as:
$\mathds{P}\Big\{x_{x_{0}\nu}(t_{k})\in\mathbb{X}_{u}~\text{for some}~k\in{[0,1,\dots\mathcal{T}]}\,\,\big|\,\,x_{0},\vartheta_{0}\Big\}$
[…]
If $\eta=0$ in (3.2c) (i.e., $\mathcal{B}$ being a nonnegative supermartingale), the results of (3.3) can be extended to infinite time horizons as
$\mathds{P}\Big\{x_{x_{0}\nu}(t_{k})\!\in\!\mathbb{X}_{u}~\text{for some}~k\in[0,\infty)\,\,\big|\,\,x_{0},\vartheta_{0}\Big\}={\epsilon}\leq\frac{\mu}{\beta}.$ (3.4)
Proof. Based on condition (3.2b), we have $\mathbb{X}_{u}\subseteq\{(x,\vartheta)\in\mathbb{X}\,\,\big|\,\,\mathcal{B}(x,\vartheta)\geq\beta\}$. Then one has
$\mathds{P}$ $\Big\{(x(t_{k}),\vartheta(t_{k}))\in\mathbb{X}_{u}\text{ for some}~0\leq k\leq\mathcal{T}\,\,\big|\,\,x_{0},\vartheta_{0}\Big\}$
$\leq\mathds{P}\Big\{\sup_{0\leq k\leq\mathcal{T}}\mathcal{B}(x(t_{k}),\vartheta(t_{k}))\geq\beta\,\,\big|\,\,x_{0},\vartheta_{0}\Big\}.$ (3.5)
[…]
Searching for control barrier certificates for large-scale dynamical systems is in general computationally very expensive. This has been the main motivation in this work to propose A-CSBC for each $\mathcal{A}_{i}(\Phi_{i})$ instead of searching for A-CBC over the interconnected A-SHS, which is not possible in practice. In the next section, we propose a compositional approach based on small-gain reasoning to construct an A-CBC of the interconnected A-SHS based on A-CSBC of individual subsystems.
4 Compositional Construction of A-CBC
Here, we first introduce a formal definition for the interconnected A-SHS similar to Definition 2.6.
Definition 4.1.
Given $N\in\mathbb{N}^{+}$ A-SHSs $\mathcal{A}_{i}(\Phi_{i})=(\mathbb{X}_{i},$ $\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},\mathbb{H}_{i})$ with their disturbances partitioned similar to (2.6), the interconnection of $\mathcal{A}_{i}(\Phi_{i})$, $\forall i\in\{1,\ldots,N\}$, is $\mathcal{A}(\Phi)=(\mathbb{X},\mathbb{U},\sigma,\rho,\varsigma,\mathbb{F},\mathbb{Y},\mathbb{H})$, denoted by $\mathcal{I}(\mathcal{A}_{1}(\Phi_{1}),\ldots,\mathcal{A}_{N}(\Phi_{N}))$, such that $\mathbb{X}\!:=\!\prod_{i=1}^{N}\mathbb{X}_{i},\mathbb{U}=\Pi_{i=1}^{N}\mathbb{U}_{i},\sigma\!:=\!\mathsf{blkdiag}(\sigma_{1}(x_{1}),\ldots,\sigma_{N}(x_{N})),$ $\rho:=\mathsf{blkdiag}(\rho_{1}(x_{1}),\ldots,\rho_{N}(x_{N})),\varsigma:=[\varsigma_{1};\dots;\varsigma_{N}],\mathbb{Y}:=\prod_{i=1}^{N}\mathbb{Y}_{i},\mathbb{H}=\prod_{i=1}^{N}\mathbb{H}_{i}$, and the map $\mathbb{F}=[\mathbb{F}_{1};\dots;\mathbb{F}_{N}]$ is defined by $(x^{\prime},\vartheta^{\prime})\!=\!\mathbb{F}((x,\vartheta),\nu,\sigma,\rho,\varsigma)$ if and only if, for any $i\in\{1,\dots,N\}$:
  1. (i)
    Flow case: $0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1$, $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, and $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
  2. (ii)
    Jump case: $\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}}$, $x_{i}^{\prime}=f_{2_{i}}(x_{i},\nu_{i},w_{i},\varsigma_{i})$, and $\vartheta_{i}^{\prime}=0$;
where $x=[x_{1};\dots;x_{N}],\vartheta=[\vartheta_{1};\dots;\vartheta_{N}],\varepsilon_{1}=[\varepsilon_{1_{i}};\dots;\varepsilon_{1_{N}}],\varepsilon_{2}=[\varepsilon_{2_{i}};\dots;\varepsilon_{2_{N}}]$ and subjected to the following interconnection constraint:
$\forall i,j\in\{1,\dots,N\},i\neq j\!:~~w_{ij}=x_{j},~~\mathbb{X}_{j}\subseteq\mathds{W}_{ij},$
where $\mathds{W}_{i}:=\prod_{j}\mathds{W}_{ij}$.
$\mathds{E}$ $\Big[\mathcal{B}(x^{\prime},\vartheta^{\prime})\,\big|\,x,\nu,\vartheta\Big]=\mathds{E}\Big[\sum_{i=1}^{N}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\,\big|\,x,\nu,\vartheta\Big]=\sum_{i=1}^{N}\xi_{i}\mathds{E}\Big[\mathcal{B}_{i}(x_{i},\vartheta_{i})\,\big|\,x,\nu,\vartheta\Big]=\sum_{i=1}^{N}\xi_{i}\mathds{E}\Big[\mathcal{B}_{i}(x_{i},\vartheta_{i})\,\big|\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big]$
$\leq\sum_{i=1}^{N}\xi_{i}\big(\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\varphi_{i}\|w_{i}\|^{2}+\eta_{i}\big)\leq\sum_{i=1}^{N}\xi_{i}\big(\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\varphi_{i}\sum_{j=1,i\neq{j}}^{N}\|w_{ij}\|^{2}+\eta_{i}\big)$
$=\sum_{i=1}^{N}\xi_{i}\big(\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\varphi_{i}\sum_{j=1,i\neq{j}}^{N}\|x_{j}\|^{2}+\eta_{i}\big)\leq\sum_{i=1}^{N}\xi_{i}\big(\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\!\!\!\!\!\sum_{j=1,i\neq{j}}^{N}\frac{\varphi_{i}}{\kappa_{j}}\mathcal{B}_{j}(x_{j},\vartheta_{j})+\eta_{i}\big)$
$=\sum_{i=1}^{N}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\sum_{i=1}^{N}\xi_{i}\big(-\theta_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\sum_{j=1,i\neq{j}}^{N}\psi_{ij}\mathcal{B}_{j}(x_{j},\vartheta_{j})+\eta_{i}\big)$
$=\mathcal{B}(x,\vartheta)+\xi^{\top}(-\Theta+\Psi)\underbrace{\big[\mathcal{B}_{1}(x_{1},\vartheta_{1});\ldots;\mathcal{B}_{N}(x_{N},\vartheta_{N})\big]}_{\tilde{\mathcal{B}}(x,\vartheta)}+\sum_{i=1}^{N}\xi_{i}\eta_{i}\leq\underbrace{(1+\Gamma)}_{\gamma}\mathcal{B}(x,\vartheta)+\eta.$ (4.3)
In the next theorem, we propose the main compositionality result of the work, using which one can compositionally construct an A-CBC for the interconnected A-SHS based on A-CSBCs of individual subsystems. To do so, we first define $\Theta:=\mathsf{diag}(\theta_{1},\ldots,\theta_{N})$ with $\theta_{i}=1-\gamma_{i}$, and $\Psi:=\{\psi_{ij}\}$ with $\psi_{ij}=\frac{\varphi_{i}}{\kappa_{j}}$, where $\psi_{ii}=0$, $\forall i\in\{1,\cdots,N\}$.
Theorem 4.2.
Consider an interconnected A-SHS $\mathcal{I}(\mathcal{A}_{1}(\Phi_{1}),\ldots,\mathcal{A}_{N}(\Phi_{N}))$ induced by $N\in\mathbb{N}^{+}$ A-SHSs $\mathcal{A}_{i}(\Phi_{i})$. Suppose that each $\mathcal{A}_{i}(\Phi_{i})$ admits an A-CSBC $\mathcal{B}_{i}$, as defined in Definition 3.1. For a vector $\xi$ with $\xi_{i}>0$, and a scalar value $\Gamma\in(-1,0)$, where $\max_{1\leq i\leq N}\Gamma_{i}<\Gamma<0$, $\forall i\in\{1,\dots,N\}$, if
$\xi^{\top}(-\Theta+\Psi)$ $:=\big[\xi_{1}\Gamma_{1};\dots;\xi_{N}\Gamma_{N}\big]^{\top}<0,$ (4.1a)
$\sum_{i=1}^{N}\xi_{i}\beta_{i}$ $>\sum_{i=1}^{N}\xi_{i}\mu_{i},$ (4.1b)
then
$\mathcal{B}(x,\vartheta):=\sum_{i=1}^{N}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})$ (4.2)
is an A-CBC for the interconnected A-SHS with the initial and unsafe sets $\mathbb{X}_{0}:=\prod_{i=1}^{N}\mathbb{X}_{0_{i}}$, $\mathbb{X}_{u}:=\prod_{i=1}^{N}\mathbb{X}_{u_{i}}$, with
$\beta:=\sum_{i=1}^{N}\xi_{i}\beta_{i},\;\mu:=\sum_{i=1}^{N}\xi_{i}\mu_{i},\;\eta:=\sum_{i=1}^{N}\xi_{i}\eta_{i},\;\gamma:=1+\Gamma.$
Proof. We first show that conditions (3.2a)-(3.2b) hold true. For any $(x,\vartheta)\in\mathbb{X}_{0}=\prod_{i=0}^{N}\mathbb{X}_{0_{i}}$ and from (3.1b), one has
$\mathcal{B}(x,\vartheta)=\sum_{i=1}^{N}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\leq\sum_{i=1}^{N}\xi_{i}\mu_{i}=\mu.$
Moreover, for any $(x,\vartheta)\in\mathbb{X}_{u}=\prod_{i=1}^{N}\mathbb{X}_{u_{i}}$ and from (3.1c), we have
$\mathcal{B}(x,\vartheta)=\sum_{i=1}^{N}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\geq\sum_{i=1}^{N}\xi_{i}\beta_{i}=\beta,$
implying that conditions (3.2a)-(3.2b) are satisfied with $\mu=\sum_{i=1}^{N}\xi_{i}\mu_{i}$ and $\beta=\sum_{i=1}^{N}\xi_{i}\beta_{i}$. Note that $\beta>\mu$ according to (4.1b).
Now we proceed with showing that condition (3.2c) holds, as well. By leveraging conditions (3.1a),(4.1a), one can acquire the chain of inequalities in (4.3). By defining
$\gamma s$ $:=\max\Big\{s+\xi^{\top}(-\Theta+\Psi)\tilde{\mathcal{B}}(x,\vartheta)\,\big|\,\xi^{\top}\tilde{\mathcal{B}}(x,\vartheta)=s\Big\},$
$\eta$ $:=\sum_{i=1}^{N}\xi_{i}\eta_{i},$
with $\tilde{\mathcal{B}}(x,\vartheta)=\big[\mathcal{B}_{1}(x_{1},\vartheta_{1});\ldots;\mathcal{B}_{N}(x_{N},\vartheta_{N})\big]$, condition (3.2c) is also met. The only remaining part is to show that $\gamma=1+\Gamma$ and $0<\gamma<1$. Given that $\xi^{\top}(-\Theta+\Psi):=\big[\xi_{1}\Gamma_{1};\dots;\xi_{N}\Gamma_{N}\big]^{\top}<0$, and $\max_{1\leq i\leq N}\Gamma_{i}<\Gamma<0$ with $\Gamma\in(-1,0)$, we have
$\gamma s$ $=s+\xi^{\top}(-\Theta+\Psi)\tilde{\mathcal{B}}(x,\vartheta)$
$=s+\big[\xi_{1}\Gamma_{1};\dots;\xi_{N}\Gamma_{N}\big]^{\top}\big[\mathcal{B}_{1}(x_{1},\vartheta_{1});\ldots;$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\mathcal{B}_{N}(x_{N},\vartheta_{N})\big]$
$=s+\xi_{1}\Gamma_{1}\mathcal{B}_{1}(x_{1},\vartheta_{1})+\dots+\xi_{N}\Gamma_{N}\mathcal{B}_{N}(x_{N},\vartheta_{N})$
$\leq s+\Gamma\big(\xi_{1}\mathcal{B}_{1}(x_{1},\vartheta_{1})+\dots+\xi_{N}\mathcal{B}_{N}(x_{N},\vartheta_{N})\big)$
$=s+\Gamma s=(1+\Gamma)s.$
[…]
Remark 4.3.
As the interconnection topology becomes denser, each subsystem is influenced by more neighbors, leading to a stronger aggregate disturbance $w_{i}$. To conservatively account for this in the analysis, the associated disturbance gain $\varphi_{i}$ may also increase, reflecting higher sensitivity to inter-subsystem coupling. Consequently, the compositional condition (4.1a), which relies on bounding this influence, becomes more difficult to satisfy, consistent with small-gain reasoning in stability analysis (see, e.g., (Dashkovskiy et al., 2007)).
As demonstrated in Theorem 4.2, under the compositionality conditions of (4.1a)-(4.1b), an A-CBC for the interconnected A-SHS can be compositionally constructed based on A-CSBCs of individual subsystems. In the next section, we provide the sufficient conditions for constructing the A-CSBC for the augmented SHS.
5 Construction of A-CSBC
In this section, we establish conditions on the SHS $\Phi_{i}$ that allow for the construction of an A-CSBC for $\mathcal{A}_{i}(\Phi_{i})$. To do so, we first define a notion of control barrier certificates (CBC) over original $\Phi_{i}$ without considering the sampling or jump constraints (Nejati et al., 2020; Anand et al., 2022). This intermediate formulation serves as a basis for systematically extending the certificates to the augmented setting to capture hybrid timing behavior.
Definition 5.1.
Given an SHS $\Phi_{i}$ with initial and unsafe sets $X_{0_{i}},X_{u_{i}}\subseteq X_{i}$, a function $\bar{\mathcal{B}}_{i}:X_{i}\rightarrow\mathbb{R}_{0}^{+}$ is called a control barrier certificate (CBC) for $\Phi_{i}$ if there exist $\gamma_{1_{i}}\in{\mathbb{R}},\bar{\kappa}_{i},\gamma_{2_{i}}\in{\mathbb{R}}^{+}$, $\bar{\mu}_{i},\bar{\beta}_{i},\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}},\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\in{\mathbb{R}}_{0}^{+}$, such that
$\bar{\mathcal{B}}_{i}(x_{i})\geq\bar{\kappa}_{i}\|x_{i}\|^{2},\quad\quad~~~~~\!\forall x_{i}\in X_{i},$ (5.1a)
$\bar{\mathcal{B}}_{i}(x_{i})\leq\bar{\mu}_{i},\quad\quad\quad\quad\quad\quad\!\!\forall x_{i}\in X_{0_{i}},$ (5.1b)
$\bar{\mathcal{B}}_{i}(x_{i})\geq\bar{\beta}_{i},\quad\quad\quad\quad\quad\quad\!\forall x_{i}\in X_{u_{i}},$ (5.1c)
  • $\forall x_{i}\in X_{i}$, $\exists\nu_{i}\in U_{i}$ such that $\forall w_{i}\in W_{i}$,
$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})\leq-\gamma_{1_{i}}\bar{\mathcal{B}}_{i}(x_{i})+\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\bar{\eta}_{1_{i}},$ (5.1d)
where $\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})$ is the infinitesimal generator acting on $\bar{\mathcal{B}}_{i}(x_{i})$, defined as (Oksendal, 2013)
$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})$ $=\frac{\partial\bar{\mathcal{B}}_{i}(x_{i})}{\partial x_{i}}\!f_{1_{i}}\!(x_{i},\!\nu_{i},\!w_{i})\!$
$~~~+\frac{1}{2}\!\textsf{Tr}~\!(\sigma_{i}(x_{i})\sigma_{i}(x_{i})\!^{\top}\frac{\partial^{2}{\bar{\mathcal{B}}_{i}(x_{i})}}{\partial x_{i}\partial x_{j}})$
$~~~+\sum_{j=1}^{\textsf{r}_{i}}\lambda_{j_{i}}(\,\bar{\mathcal{B}}_{i}(x_{i}+\rho_{i}(x_{i})\textsf{e}_{j_{i}}^{\textsf{r}_{i}})-\bar{\mathcal{B}}_{i}(x_{i})),$
where $\lambda_{j}$ is the rate of Poisson processes and $\textsf{e}_{j}^{\textsf{r}}$ is an r-dimensional vector with $1$ on the $j$-th entry and 0 elsewhere;
  • $\forall x_{i}\in X_{i}$, $\exists\nu_{i}\in U_{i}$ such that $\forall w_{i}\in W_{i}$,
$\mathds{E}\Big[\bar{\mathcal{B}}_{i}(f_{2_{i}}(x_{i},\!\nu_{i},\!w_{i})\!)\big|x_{i},\!\nu_{i},\!w_{i}\Big]\!\!\leq\!\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\!\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}.$ (5.1e)
The above definition introduces a stochastic CBC $\bar{\mathcal{B}}_{i}$ for the SHS $\Phi_{i}$, analyzing the system behavior. For instance, with $\gamma_{1_{i}}>0$ and $0<\gamma_{2_{i}}<1$, it ensures that the expected certificate value decreases (or remains bounded) during both flow and jump phases (conditions (5.1d) and (5.1e), respectively). It is worthwhile noting that a thorough analysis is conducted for other bounds of $\gamma_{1_{i}}$ and $\gamma_{2_{i}}$ in Theorem 5.4 (cf. (5.8)). The terms $\bar{\varphi}_{i_{1}}$ and $\bar{\varphi}_{i_{2}}$ capture the effect of the adversarial input, enabling this certificate to quantify robustness against such disturbances, while for instance $\gamma_{1_{i}}>0$ and $0<\gamma_{2_{i}}<1$ characterize exponential decay or boundedness of the certificate in each mode.
Remark 5.2.
Definition 5.1 formulates CBCs tailored to SHSs that exhibit both continuous-time and discrete-time (reset) dynamics, as defined in Definition 2.1. In this setting, in the case that $\gamma_{1_{i}}>0$ and $0<\gamma_{2_{i}}<1$, condition (5.1d) ensures the desired decrease of the certificate along the continuous-time evolution, using the infinitesimal generator commonly employed in continuous-time stochastic analysis (Nejati et al., 2020). In contrast, condition (5.1e) governs the discrete-time updates and captures the expected decrease across discrete transitions, consistent with the discrete-time stochastic setting discussed in (Anand et al., 2022). Together, these conditions jointly ensure the safety of the SHS $\Phi_{i}$ under both modes of evolution.
Before offering the main result of this section, we also raise the following lemma to quantify an upper bound on the evolution of CBC $\bar{\mathcal{B}}_{i}$.
Lemma 5.3.
Given an SHS $\Phi_{i}=(X_{i},U_{i},\mathcal{U}_{i},W_{i},\mathcal{W}_{i},\sigma_{i},$ $\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$, suppose (5.1d) holds. Then for any $x_{i}\in X_{i},\nu_{i}\in\mathcal{U}_{\tau_{i}},w_{i}\in W_{i}$, and for any two consecutive time instances $(t_{k},t_{k+1})$, one has
$\mathds{E}$ $\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
$\leq e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k})\!)$
$~~~+e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}}).$ (5.2)
Proof. From Dynkin’s formula (Dynkin, 1965), we have:
$\mathds{E}$ $\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
$=\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))\!+\!\mathds{E}\Big[\int_{t_{k}}^{t_{k+1}}\!\!\!\!\!\!\!\mathcal{L}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))\mathsf{d}t\Big]\!.$
From condition (5.1d), one has
$\mathds{E}$ $\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big|\,x_{i},\nu_{i},w_{i}\Big]\leq\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))$
$+\mathds{E}\Big[\int_{t_{k}}^{t_{k+1}}\big(-\gamma_{1_{i}}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))+\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\bar{\eta}_{1_{i}}\big)\mathsf{d}t$
$=\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))\!+\!(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}})$
$~~~-\int_{t_{k}}^{t_{k+1}}\!\!\gamma_{1_{i}}\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))\Big]\mathsf{d}t.$
By employing Gronwall inequality (Gronwall, 1919), we have
$\mathds{E}$ $\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
$\leq e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))$
$+e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}}),$
[…]
Under Definition 5.1 and Lemma 5.3, we propose the next theorem to construct an A-CSBC for $\mathcal{A}_{i}(\Phi_{i})$ based on the CBC $\bar{\mathcal{B}}_{i}$.
Theorem 5.4.
Given an SHS $\Phi_{i}=(X_{i},U_{i},\mathcal{U}_{i},W_{i},\mathcal{W}_{i},$ $\sigma_{i},\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$ with its associated A-SHS $\mathcal{A}_{i}(\Phi_{i})=(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},\mathbb{H}_{i})$, suppose $\bar{\mathcal{B}}_{i}$ is a CBC for $\Phi_{i}$. If
$\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\vartheta_{i}<0,\quad\forall\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\},$ (5.3)
then
$\mathcal{B}_{i}(x_{i},\vartheta_{i})=\varpi_{i}~\!\bar{\mathcal{B}}_{i}(x_{i})$ (5.4)
is an A-CSBC for $\mathcal{A}_{i}(\Phi_{i})$ with
$\varpi_{i}:=\left\{\hskip-1.42262pt\begin{array}[]{lr}1,\quad\quad\quad\,~~~~\text{if}~\gamma_{1_{i}}>0~\&~0<\gamma_{2_{i}}<1,\\ e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}},\quad\text{if}~\gamma_{1_{i}}>0~\&~\gamma_{2_{i}}\geq 1,\\ \gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}},\quad\quad~~~~\!\text{if}~\gamma_{1_{i}}\leq 0~\&~0<\gamma_{2_{i}}<1,\end{array}\right$. (5.8)
for some $0<\alpha_{1_{i}}<1$ and $\alpha_{2_{i}}>\varepsilon_{2_{i}}$.
Proof. By employing (5.1a), for any $(x_{i},\vartheta_{i})\in\mathbb{X}_{i}$, one has
$\mathcal{B}_{i}(x_{i},\vartheta_{i})=\varpi_{i}~\!\bar{\mathcal{B}}_{i}(x_{i})\geq\varpi_{i}\bar{\kappa}_{i}\|x_{i}\|^{2}.$
Then condition (3.1a) holds with $\kappa_{i}=\varpi_{i}\bar{\kappa}_{i}$. By leveraging (5.1b), for any $(x_{i},\vartheta_{i})\in\mathbb{X}_{0_{i}}$, one has
$\mathcal{B}_{i}(x_{i},\vartheta_{i})=\varpi_{i}~\!\bar{\mathcal{B}}_{i}(x_{i})\leq\varpi_{i}~\!\bar{\mu}_{i},$
and similarly by utilizing (5.1c) for any $(x_{i},\vartheta_{i})\in\mathbb{X}_{u_{i}}$, one has
$\mathcal{B}_{i}(x_{i},\vartheta_{i})=\varpi_{i}~\!\bar{\mathcal{B}}_{i}(x_{i})\geq\varpi_{i}~\!\bar{\beta}_{i},$
satisfying conditions (3.1b) and (3.1c) with $\mu_{i}=\varpi_{i}\bar{\mu}_{i}$ and $\beta_{i}=\varpi_{i}\bar{\beta}_{i}$.
Now we proceed with showing condition (3.1d), as well. From (5.2) with $t_{k+1}=\tau,t_{k}=0$, one gets
$\mathds{E}$ $\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
$\leq e^{-\gamma_{1_{i}}\tau}\big(\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(0))\!+\!\tau(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}})\big),$
$=e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})+e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}.$
In order to show that $\mathcal{B}_{i}$ defined in (5.4) fulfills condition (3.1d), we consider the two cases in Definition 2.8 and the three bounds for different values of $\gamma_{1_{i}}$ and $\gamma_{2_{i}}$ in (5.8) as follows:
  • Bound 1 ($\gamma_{1_{i}}>0~\&~0<\gamma_{2_{i}}<1$):
    1. (i)
      Flow case $(0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1,\vartheta_{i}^{\prime}=\vartheta_{i}+1)$:
      $\mathds{E}\Big[\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big|\,x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big]\!\!=\!\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big|\,x_{i},\!\nu_{i},\!w_{i}\Big]$
      $\leq e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})\!+\!e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}\!+\!e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~~~+\!e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}\!+\!\|w_{i}\|^{2})$
      $~~~+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}$
      $~~~+e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}.$
    2. (ii)
      Jump case $(\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}},\vartheta_{i}^{\prime}=0)$:
      $\mathds{E}$ $\Big[\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big|\,x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big]\!=\!\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big|\,x_{i},\!\nu_{i},\!w_{i}\Big]$
      $\leq\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
      $=\gamma_{2_{i}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}+\bar{\eta}_{2_{i}}.$
      By defining $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau},\gamma_{2_{i}}\}$, $\varphi_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}$ and $\eta_{i}\!=\!\max\{e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\\ \|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}$, we have
      $\mathds{E}\Big[\mathcal{B}_{i}(x_{i}^{\prime},\!\vartheta_{i}^{\prime})\big|x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big]$ $\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})$
      $~~~+\varphi_{i}\|w_{i}\|^{2}\!+\!\eta_{i}.$
  • Bound 2 ($\gamma_{1_{i}}>0~\&~\gamma_{2_{i}}\geq 1$):
    1. (i)
      Flow case $(0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1,\vartheta_{i}^{\prime}=\vartheta_{i}+1)$:
      $\mathds{E}$ $\Big[\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big]$
      $=e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}^{\prime}}\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
      $=e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
      $\leq e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}\big(e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})$
      $~~~+e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}\big)$
      $=e^{-\gamma_{1_{i}}\tau}\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}}e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}}\bar{\mathcal{B}}_{i}(x_{i})$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!$
      $~~~+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})$
      $~~~+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})$
      $~~~+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}+\|w_{i}\|^{2})$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})$
      $+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}$
      $+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}.$
    2. (ii)
      Jump case $(\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}},\vartheta_{i}^{\prime}=0)$:
      $\mathds{E}$ $\Big[\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big]$
      $=e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}^{\prime}}\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
      $=\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
      $\leq\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
      $=\frac{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}}}{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}}}\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}}\gamma_{2_{i}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}.$
      By defining
      $\gamma_{i}$ $=\max\{e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})},e^{-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}}\gamma_{2_{i}}\},$
      $\varphi_{i}$ $=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\},$
      and
      $\eta_{i}$ $=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\},$
      we have
      $\mathds{E}\Big[\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\big|x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big]$ $\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!$
      $~~~+\!\varphi_{i}\|w_{i}\|^{2}\!\!+\!\eta_{i}.$
  • Bound 3 ($\gamma_{1_{i}}\leq 0~\&~0<\gamma_{2_{i}}<1$):
    1. (i)
      Flow case $(0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1,\vartheta_{i}^{\prime}=\vartheta_{i}+1)$:
      $\mathds{E}\Big[\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big]$
      $=\gamma_{2_{i}}^{\frac{\vartheta_{i}^{\prime}}{\alpha_{2_{i}}}}\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
      $=\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
      $\leq\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}\Big(e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})+e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $\quad+e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}\Big)$
      $=e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}}\bar{\mathcal{B}}_{i}(x_{i})$
      $\quad+\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}\!+\!\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}\!e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})$
      $\quad+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})$
      $\quad+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}+\|w_{i}\|^{2})$
      $\quad+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}$
      $\quad+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}.$
    2. (ii)
      Jump case $(\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}},\vartheta_{i}^{\prime}=0)$:
      $\mathds{E}$ $\Big[\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big]$
      $=\gamma_{2_{i}}^{\frac{\vartheta_{i}^{\prime}}{\alpha_{2_{i}}}}\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
      $=\!\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\big|x_{i},\!\nu_{i},\!w_{i}\Big]$
      $\leq\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
      $=\frac{\gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}}}{\gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}}}\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
      $\leq\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}.$
      By defining
      $\gamma_{i}$ $=\max\{e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}},\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}\},$
      $\varphi_{i}$ $=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\},$
      and
      $\eta_{i}=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}$ $\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\},$
      we have
      $\mathds{E}\Big[\mathcal{B}_{i}(x_{i}^{\prime},\!\vartheta_{i}^{\prime})\big|x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big]$ $\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})$
      $~~~+\varphi_{i}\|w_{i}\|^{2}+\eta_{i}.$
To complete the proof, we proceed with showing that $0<\gamma_{i}<1$ for all Bounds 1-3. One can readily verify that
$0<\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau},\gamma_{2_{i}}\}<1,$
for Bound 1. In Bound 2, since $0<\alpha_{1_{i}}<1$ and $\gamma_{1_{i}}>0$, then we have $e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}<1$. In addition,
$e^{-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}}\gamma_{2_{i}}<1\Leftrightarrow\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}<0.$
Since $\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\vartheta_{i}<0$ for any $\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\}$ and by leveraging the continuity property of real numbers, one can always find some $0<\alpha_{1_{i}}<1$ such that $\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}<0$. Hence, $0<\gamma_{i}<1$. Similarly, in Bound 3, one has $\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}<1$ since $\alpha_{2_{i}}>\varepsilon_{2_{i}}$ and $0<\gamma_{2_{i}}<1$. Additionally,
$e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}<1\Leftrightarrow\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\alpha_{2_{i}}<0.$
[…]
Theorem 5.4 establishes that the A-CSBC conditions are fulfilled under the transformation $\mathcal{B}_{i}(x_{i},\vartheta_{i})=\varpi_{i}\bar{\mathcal{B}}_{i}(x_{i})$, provided that condition (5.3) holds. The proposed approach systematically addresses all hybrid transitions—both flows and jumps—by categorizing them into three distinct cases, corresponding to the different decay or growth patterns exhibited by the original CBC.
Remark 5.5.
The introduction of multiple definitions of control barrier certificates in this work serves distinct but complementary purposes within our framework. Definition 3.1 introduces A-CSBCs tailored to the A-SHS model defined in Definition 2.8. While A-CSBCs are necessary to ensure compositional safety using subsystems, their design is complicated by the inclusion of the auxiliary counter $\vartheta_{i}$. In contrast, Definition 5.1 presents a conventional CBC formulation (excluding the effect of $\vartheta_{i}$) for SHS by combining the results of stochastic non-hybrid continuous- and discrete-time systems, which serves as the foundational tool for safety analysis. To bridge the gap between these two definitions, Theorem 5.4 establishes a systematic method to construct A-CSBCs from conventional CBCs by introducing an auxiliary parameter $\varpi_{i}$.
Remark 5.6.
The core challenge in analyzing networked stochastic hybrid systems lies not only in their high dimensionality but also in the hybrid coupling between subsystems, captured by the adversarial inputs $w_{i}$ in (2.4). This coupling significantly impacts the system’s behavior and should be explicitly considered in our formal safety analysis. To this end, our framework incorporates $w_{i}(t)$ throughout the analysis: they are integrated in our development in Lemma 5.3, their bounds are formally established in Theorem 4.2, and their impact is captured by the parameter $\varphi_{i}$, introduced in Definition 3.1. This parameter is then properly designed in Theorem 5.4—as shown in its proof—under three distinct scenarios (Bounds 1–3), that can cover the hybrid nature of dynamics.
6 Computation of CBC
Here, we employ sum-of-squares (SOS) optimization techniques (Parrilo, 2003) and search for a CBC and its corresponding controller for each $\Phi_{i}$. To do so, we assume that $f_{1_{i}}$ and $f_{2_{i}}$ are polynomial functions. The next lemma provides the SOS formulation.
Lemma 6.1.
Suppose sets $X_{0_{i}},X_{u_{i}},X_{i},U_{i},$ and $W_{i}$ can be described by $X_{0_{i}}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{0_{i}}(x_{i})\geq 0\}$, $X_{u_{i}}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{u_{i}}(x_{i})\geq 0\}$, $X_{i}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{i}(x_{i})\geq 0\}$, $U_{i}=\{U_{i}\subseteq{\mathbb{R}}^{m_{i}}\mid g_{\nu_{i}}(\nu_{i})\geq 0\}$, and $W_{i}=\{W_{i}\subseteq{\mathbb{R}}^{p_{i}}\mid g_{w_{i}}(w_{i})\geq 0\}$. Let there exist an SOS polynomial $\bar{\mathcal{B}}_{i}(x_{i})$, constants $\gamma_{1_{i}}\in{\mathbb{R}},\bar{\kappa}_{i},\gamma_{2_{i}}\in{\mathbb{R}}^{+}$, $\bar{\mu}_{i},\bar{\beta}_{i},\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}},\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\in{\mathbb{R}}_{0}^{+}$, polynomial $l_{\nu_{j_{i}}}(x_{i})$, corresponding to the $j^{\text{th}}$ input in $\nu_{i}=(\nu_{1_{i}},\nu_{2_{i}},\ldots,\nu_{m_{i}})\in U_{i}\subseteq{\mathbb{R}}^{m_{i}}$. Suppose there exist vectors of SOS polynomials $l_{0_{i}}(x_{i})$, $l_{u_{i}}(x_{i})$, $\tilde{l}_{i}(x_{i})$, $l_{i}(x_{i},\nu_{i},w_{i}),\hat{l}_{i}(x_{i},\nu_{i},w_{i}),l_{\nu_{i}}(x_{i},\nu_{i},w_{i}),l_{w_{i}}(x_{i},\nu_{i},w_{i}),\\ \hat{l}_{w_{i}}(x_{i},\nu_{i},w_{i})$ of appropriate dimensions such that the following expressions are SOS polynomials:
$\bar{\mathcal{B}_{i}}(x_{i})-\tilde{l}_{i}^{\top}(x_{i})g_{i}(x_{i})-\bar{\kappa}_{i}x_{i}^{\top}x_{i},$ (6.1a)
- $\bar{\mathcal{B}_{i}}(x_{i})-l_{0_{i}}^{\top}(x_{i})g_{0_{i}}(x_{i})+\bar{\mu}_{i},$ (6.1b)
$\bar{\mathcal{B}}_{i}(x_{i})\!-\!l_{u_{i}}^{\top}(x_{i})g_{u_{i}}(x_{i})\!-\!\bar{\beta}_{i},$ (6.1c)
- $\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})\!-\!\gamma_{1_{i}}\!(\bar{\mathcal{B}}_{i}(x_{i})\!)\!+\!\bar{\varphi}_{i_{1}}w_{i}^{\top}w_{i}\!+\!\bar{\eta}_{1_{i}}\!-\!\!\!\sum_{j=1}^{m_{i}}(\nu_{j_{i}}\!-\!l_{\nu_{j_{i}}}(x_{i})\!)$
$-l_{i}^{\top}(x_{i},\nu_{i},w_{i})g_{i}(x_{i})\!-\!l_{\nu_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{\nu_{i}}(\nu_{i})$
$-l_{w_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{w_{i}}(w_{i}),$ (6.1d)
- $\mathds{E}\Big[\bar{\mathcal{B}}_{i}(f_{2_{i}}(x_{i},\nu_{i},w_{i})\!)\big|x_{i},\!\nu_{i},\!w_{i}\Big]\!+\!\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}w_{i}^{\top}w_{i}$
$+\bar{\eta}_{2_{i}}\!-\!\sum_{j=1}^{m_{i}}(\nu_{j_{i}}\!-\!{l_{\nu_{j_{i}}}(x_{i})}\!)-\hat{l}_{i}^{\top}(x_{i},\nu_{i},w_{i})g_{i}(x_{i})$
$-{l_{\nu_{i}}^{\top}(x_{i},\nu_{i},w_{i})}g_{\nu_{i}}(\nu_{i})-\hat{l}_{w_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{w_{i}}(w_{i}).$ (6.1e)
Then $\bar{\mathcal{B}}_{i}(x_{i})$ fulfills conditions (5.1a)-(5.1e) in Definition 5.1. Furthermore, $\nu_{i}=[l_{\nu_{1_{i}}}(x_{i});\dots;l_{\nu_{m_{i}}}(x_{i})]$ is the corresponding safety controller.
7 Case Study
Here, we demonstrate the effectiveness of our approach by applying it to an interconnected stochastic hybrid system, composed of 1000 nonlinear subsystems with two distinct topologies—one with a ring topology and the other with a 3-in-3-out interconnection topology—presented in Subsections 7.1 and 7.2, respectively.
7.1 Ring Interconnection Topology
Here, we demonstrate the effectiveness of our proposed results by applying them to an interconnected SHS composed of $1000$ nonlinear subsystems. The model of the interconnected SHS $\Phi$ is given as
$\Phi\!:\left\{\hskip-4.2679pt\begin{array}[]{rl}\mathsf{d}x(t)\!=\!\!&(A_{1}x^{\circ 3}(t)\!+\!B_{1}\nu(t))\mathsf{d}t+E\mathsf{d}\mathbb{W}_{t}+F\mathsf{d}\mathbb{P}_{t},\\ &~\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad~t\in\mathbb{R}_{0}^{+}\backslash\Delta,\\ x(t)\!=\!\!&A_{2}x^{\circ 3}(t^{-})\!+\!B_{2}\nu(t)\!+\!G\varsigma(t),\quad~~~\!t\in\Delta,\end{array}\right$. (7.4)
where $A_{1},A_{2}$ are matrices with diagonal elements $\bar{a}_{1_{ii}}=a_{1},\bar{a}_{2_{ii}}=a_{2}$, off-diagonal elements $\bar{a}_{1_{i,i-1}}=\bar{a}_{1_{1,N}}=d_{1}$, $\bar{a}_{2_{i,i-1}}=\bar{a}_{2_{1,N}}=d_{2}$, $i\in\{2,\ldots,N\}$, and all other elements are identically zero. Moreover, $B_{1}=\mathsf{diag}(b_{1})$, $B_{2}=\mathsf{diag}(b_{2})$, $E=0.6\mathds{I}_{N},F=0.5\mathds{I}_{N},G=0.5\mathds{I}_{N}$, $x=[x_{1};\ldots;x_{N}]$, $\nu(t)=[\nu_{1}(t);\ldots;{\nu_{N}(t)}]$. Now by considering individual subsystems as $\Phi_{i}$ as
$\Phi_{i}\!\!:\!\left\{\hskip-4.2679pt\begin{array}[]{rl}\mathsf{d}x_{i}(t)\!=\!\!&(a_{1}x_{i}^{3}(t)\!+\!b_{1}\nu_{i}(t)\!+\!d_{1}w_{i}(t))\mathsf{d}t+0.6\mathsf{d}\mathbb{W}_{t}\\ \quad\quad&+~0.5\mathsf{d}\mathbb{P}_{t},\quad\quad\quad\quad\quad\quad\quad\quad\quad t\in\mathbb{R}_{0}^{+}\backslash\Delta_{i},\\ x_{i}(t)\!=\!&a_{2}x_{i}^{3}(t^{-})\!+\!b_{2}\nu_{i}(t)\!+\!d_{2}{w_{i}(t^{-})}\!+\!0.5\varsigma_{i}(t),\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad~t\!\in\!\Delta_{i},\end{array}\right$. (7.9)
with $a_{1},b_{1},a_{2},b_{2}\in\mathbb{R}$, jump parameters $\tau=0.1$, $\varepsilon_{1_{i}}=1,\varepsilon_{2_{i}}=7$, and rates of Poisson processes as $\lambda_{i}=0.5$, one can readily verify that $\Phi=\mathcal{I}(\Phi_{1},\ldots,\Phi_{N})$ where $w_{i}(t)=x_{i-1}(t)$ (with $x_{0}=x_{N}$), i.e., ring interconnection topology.
The regions of interest are given as $X\in[0,10]^{1000},X_{0}\in[0,2]^{1000},X_{u}=[9,10]^{1000}$. The main goal is to design an A-CBC for the interconnected A-SHS and its corresponding safety controllers based on A-CSBCs of individual subsystems such that the state of the system remains in the comfort zone $0,9]^{1000}$. To do so, we first search for CBCs and accordingly design local controllers for $\Phi_{i}$. Consequently, A-CSBCs can be constructed from CBCs according to Theorem 5.4. We show our results for different values of $a_{1},b_{1},a_{2},b_{2}$.
We employ the toolbox SOSTOOLS (Papachristodoulou et al., 2013) and the SDP solver SeDuMi (Sturm, 1999) to compute CBC as described in Section 6. For $a_{1}=-0.4,b_{1}=0.5,a_{2}=0.01,b_{2}=0.06$, based on Lemma 6.1, we compute a CBC of an order $4$ as $\bar{\mathcal{B}_{i}}(x_{i})=0.0029x_{i}^{4}-0.0189^{3}+0.0458x_{i}^{2}-0.0494x_{i}+0.0369$ and corresponding controller $\nu_{i}=-0.09x_{i}+2.7$. Moreover, the corresponding constants in Definition 5.1 satisfying conditions (5.1a)-(5.1e) are quantified as $\bar{\mu}_{i}=0.0684,~\bar{\beta}_{i}=7.2219,~\bar{\kappa}_{i}=0.011$, $\bar{\eta}_{1_{i}}=0.0015,~\bar{\eta}_{2_{i}}=0.0012,~\bar{\varphi}_{1_{i}}=0.00001,~\bar{\varphi}_{2_{i}}=0.00001,~\gamma_{1_{i}}=0.01,~\gamma_{2_{i}}=0.99$. We now proceed with Theorem 5.4 to construct an A-CSBC for the augmented system using the obtained CBC. We select $\alpha_{1_{i}}=0.1,\alpha_{2_{i}}=8$. Since $\gamma_{1_{i}}=0.01>0$ and $0<\gamma_{2_{i}}=0.99<1$, the results of Bound 1 are useful. One can readily verify that condition (5.3) is also met for all $\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\}$. Then one can conclude that $\mathcal{B}_{i}(x_{i},\vartheta_{i}):=\bar{\mathcal{B}}_{i}(x_{i})$ is a CSBC for the augmented SHS with $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau},\gamma_{2_{i}}\}=0.999$, $\varphi_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}=0.00001$ and $\eta_{i}\!=\!\max\{e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}=0.0012$. We now proceed to construct an A-CBC based on A-CSBCs of individual subsystems according to the compositionality Theorem 4.2. By selecting $\xi_{i}=0.001$$\forall i\in\{1,\dots,1000\}$, one has $\xi_{i}\Gamma_{i}=-9.0869\times\!10^{-6}\!<\!0,$ implying the satisfaction of condition (4.1a). Moreover, we compute $\beta=\sum_{i=1}^{1000}\xi_{i}\beta_{i}=7.2219$, and $\mu=\sum_{i=1}^{1000}\xi_{i}\mu_{i}=0.0684,$ that imply condition (4.1b). Then, it can be concluded that, $\mathcal{B}(x,\vartheta)=\sum_{i=1}^{1000}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i}),$ satisfying conditions in (3.2) with $\eta=\sum_{i=1}^{N}\xi_{i}\eta_{i}=0.0012,$ and $\gamma=1+\Gamma=0.9909$.
Figure 1: Control barrier certificate (top) and closed-loop state trajectories (bottom) of a representative subsystem within the interconnected SHS $\Phi$ with ring topology with $10$ noise realizations for $a_{1}=-0.4,b_{1}=0.5,a_{2}=0.01,b_{2}=0.06$ (left), $a_{1}=-0.3,b_{1}=0.2,a_{2}=1.01,b_{2}=1$ (middle), $a_{1}=0.01,b_{1}=0.7,a_{2}=0.075,b_{2}=0.8$ (right). Initial and unsafe regions are depicted by green and red boxes. Moreover, $\bar{\mathcal{B}}_{i}(x_{i})=\bar{\mu}_{i}$ and $\bar{\mathcal{B}}_{i}(x_{i})=\bar{\beta}_{i}$ indicated by  and  , respectively.
For $a_{1}=-0.3,b_{1}=0.2,a_{2}=1.01,b_{2}=1$, we compute a CBC of an order $4$ as $\bar{\mathcal{B}}_{i}(x_{i})=0.0013x^{4}_{i}-0.0108x^{3}_{i}+0.0328x^{2}_{i}-0.0434x_{i}+0.0281$ and corresponding controller $\nu_{i}=-1.02x^{3}_{i}+3.05$. Furthermore, the corresponding constants satisfying conditions (5.1a)-(5.1e) are obtained as $\bar{\mu}_{i}=0.0294,~\bar{\beta}_{i}=3.0157,~\bar{\kappa}_{i}=7.0587\times 10^{-4},$ $\gamma_{1_{i}}=0.04547,\gamma_{2_{i}}=1.00001,$ $\bar{\varphi}_{1_{i}}=1.3804\times 10^{-5},\bar{\varphi}_{2_{i}}=2.443\times 10^{-6}$, $\bar{\eta}_{1_{i}}=0.0028,~\bar{\eta}_{2_{i}}=0.003$. Since $\gamma_{1_{i}}=0.04547>0$ and $\gamma_{2_{i}}=1.00001\geq 1$, Bound 2 is valid. Condition (5.3) is also met for all $\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\}$. Then $\mathcal{B}_{i}(x_{i},\vartheta_{i}):=e^{4.57\times 10^{-4}\vartheta_{i}}\bar{\mathcal{B}}_{i}(x_{i})$ is a CSBC for the augmented SHS with $\mu_{i}=0.0294$, $\beta_{i}=3.0254$, $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})},e^{-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}}\gamma_{2_{i}}\}=0.9996,$ $\varphi_{i}=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}=2.443\times 10^{-6}$, and $\eta_{i}=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}$ $e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}=0.003$. We now proceed to construct an A-CBC based on A-CSBCs of individual subsystems according to the compositionality Theorem 4.2. By choosing $\xi_{i}=0.001,$ $\forall i\in\{1,\dots,1000\}$, one has $\xi_{i}\Gamma_{i}=-6.539\times\!10^{-6}\!<\!0,$ implying the satisfaction of condition (4.1a). Moreover, we compute $\beta=\sum_{i=1}^{1000}\xi_{i}\beta_{i}=3.0254$, and $\mu=\sum_{i=1}^{1000}\xi_{i}\mu_{i}=0.0294,$ that imply the satisfaction of condition (4.1b). Hence, it can be concluded that $\mathcal{B}(x,\vartheta)=\sum_{i=1}^{1000}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})$, implying the satisfaction of conditions in (3.2) with $\eta=\sum_{i=1}^{N}\xi_{i}\eta_{i}=0.003,$ and $\gamma=1+\Gamma=0.9935.$
For $a_{1}=0.01,b_{1}=0.7,$ $a_{2}=0.075,b_{2}=0.8$, we compute a CBC of an order $4$ as $\bar{\mathcal{B}}_{i}(x_{i})=0.0035x^{4}_{i}-0.0284x^{3}_{i}+0.0814x^{2}_{i}-0.0982x_{i}+0.0692$ and its corresponding controller $\nu_{i}=-0.099x^{3}_{i}+2.5$. In addition, the corresponding constants in Definition 5.1 are acquired as $\bar{\mu}_{i}=0.0727,~\bar{\beta}_{i}=8.2123,~\bar{\kappa}_{i}=0.0024,\gamma_{1_{i}}=-0.31,~\gamma_{2_{i}}=0.77$, $\bar{\varphi}_{1_{i}}=7.5654\times 10^{-6},~\bar{\varphi}_{2_{i}}=8.1243\times 10^{-6}$, and $\bar{\eta}_{1_{i}}=0.01,~\bar{\eta}_{2_{i}}=0.01$. Since $\gamma_{1_{i}}=-0.31\leq 0$ and $0<\gamma_{2_{i}}=0.77<1$, Bound 3 is useful. Condition (5.3) is also met for all $\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\}$. Consequently, one can conclude that $\mathcal{B}_{i}(x_{i},\vartheta_{i}):=0.77^{\frac{\vartheta_{i}}{8}}\bar{\mathcal{B}_{i}}(x_{i})$ is a CSBC for the augmented SHS with $\mu_{i}=0.0578$, $\beta_{i}=7.9483$, $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}},\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}\}=0.9983$, $\varphi_{i}=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}=8.1243\times 10^{-6}$, and $\eta_{i}=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}=0.01$. We now proceed to construct an A-CBC based on A-CSBCs of individual subsystems according to the compositionality Theorem 4.2. By selecting $\xi_{i}=0.001$, $\forall i\in\{1,\dots,1000\}$, one has $\xi_{i}\Gamma_{i}=-6.6123\times\!10^{-6}\!<\!0,$ implying the satisfaction of condition (4.1a). Moreover, we compute $\beta=\sum_{i=1}^{1000}\xi_{i}\beta_{i}=7.9483$, and $\mu=\sum_{i=1}^{1000}\xi_{i}\mu_{i}=0.0578,$ that imply the fulfillment of condition (4.1b). Then, it can be concluded that $\mathcal{B}(x,\vartheta)=\sum_{i=1}^{1000}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})$, satisfying conditions in (3.2) with $\eta=\sum_{i=1}^{N}\xi_{i}\eta_{i}=0.01,$ and $\gamma=1+\Gamma=0.9934.$
By employing Theorem 3.4, we guarantee that the state trajectory of $\Phi$ starting from initial conditions inside $X_{0}=[0~2]^{1000}$ remains in the safe set $0~9]^{1000}$ during the time steps $\mathcal{T}=100$ with the probability of at least $0.9742$ (first bound), $0.8937$ (second bound), and $0.8753$ (last bound). A control barrier certificate for a representative subsystem within the interconnected SHS $\Phi$ is depicted in Fig. 1. Closed-loop state trajectories of the interconnected SHS $\Phi$ for different bounds with $10$ noise realizations are depicted in Fig. 1. Note that we have only plotted trajectories at the discrete time steps since our guarantee is provided for them. The computation of CBC and its corresponding controllers for each bound took almost $4.81$ seconds with a memory usage of $64.84~\mathsf{Mbit}$ on a MacBook Pro M2 Max chip with $32$ GB of memory.
7.2 3-in-3-out Interconnection Topology
In this subsection, we consider an interconnected network comprising 1000 nonlinear subsystems, where each subsystem is influenced by three preceding subsystems and affects three following ones. The model is the same as in (7.4), where the diagonal elements of the matrices $A_{1}$ and $A_{2}$ are given by $\bar{a}_{1_{ii}}=a_{1}$ and $\bar{a}_{2_{ii}}=a_{2}$, respectively. The off-diagonal elements are defined as $\bar{a}_{1_{i,i-1}}=\bar{a}_{1_{i,i-2}}=\bar{a}_{1_{i,i-3}}=\bar{a}_{1_{1,N}}=\bar{a}_{1_{1,N-1}}=\bar{a}_{1_{1,N-2}}=\bar{a}_{1_{2,1}}=\bar{a}_{1_{2,N}}=\bar{a}_{1_{2,N-1}}=\bar{a}_{1_{3,1}}=\bar{a}_{1_{3,2}}=\bar{a}_{1_{3,N}}=d_{1},$ $\bar{a}_{2_{i,i-1}}=\bar{a}_{2_{i,i-2}}=\bar{a}_{2_{i,i-3}}=\bar{a}_{2_{1,N}}=\bar{a}_{2_{1,N-1}}=\bar{a}_{2_{1,N-2}}=\bar{a}_{2_{2,1}}=\bar{a}_{2_{2,N}}=\bar{a}_{2_{2,N-1}}=\bar{a}_{2_{3,1}}=\bar{a}_{2_{3,2}}=\bar{a}_{2_{3,N}}=d_{2},$ for all $i\in\{4,\ldots,N\}$. Each subsystem in this network is described by
$\Phi_{i}\!\!:\!\left\{\hskip-4.2679pt\begin{array}[]{rl}\mathsf{d}x_{i}(t)\!=\!\!&(a_{1}x_{i}^{3}(t)\!+\!b_{1}\nu_{i}(t)\!+\!D_{1}w_{i}(t))\mathsf{d}t+0.6\mathsf{d}\mathbb{W}_{t}\\ \quad\quad&+~0.5\mathsf{d}\mathbb{P}_{t},\quad\quad\quad\quad\quad\quad\quad\quad\quad t\in\mathbb{R}_{0}^{+}\backslash\Delta_{i},\\ x_{i}(t)\!=\!&a_{2}x_{i}^{3}(t^{-})\!+\!b_{2}\nu_{i}(t)\!+\!D_{2}w_{i}(t^{-})\!+\!0.5\varsigma_{i}(t),\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad~t\!\in\!\Delta_{i},\end{array}\right$. (7.14)
where $D_{1}=[d_{1};d_{1};d_{1}]^{\top}$, $D_{2}=[d_{2};d_{2};d_{2}]^{\top}$, and $w_{i}=[x_{i-1};x_{i-2};x_{i-3}]^{\top}$, with the wrap-around conditions $x_{0}=x_{N}$, $x_{-1}=x_{N-1}$, and $x_{-2}=x_{N-2}$. We refer to this configuration as the 3-in-3-out interconnection topology, in which the size of $w_{i}$ is three times that in the ring topology. Hence, this represents a denser interconnection structure compared to the one presented in Subsection 7.1. For the purpose of comparing the two topologies, all system parameters are kept identical to those used in Subsection 7.1.
For $a_{1}=-0.4,b_{1}=0.5,a_{2}=0.01,b_{2}=0.06$, we synthesize a CBC and its associated safety controller as $\bar{\mathcal{B}_{i}}(x_{i})=0.0021x_{i}^{4}-0.0138^{3}+0.0339x_{i}^{2}-0.0371x_{i}+0.0235$ and $\nu_{i}=-0.08x_{i}+2.6$, respectively. The parameters satisfying conditions (5.1a)-(5.1e) in Definition 5.1 are designed as $\bar{\mu}_{i}=0.049,~\bar{\beta}_{i}=5.0557,~\bar{\kappa}_{i}=0.0105$, $\bar{\eta}_{1_{i}}=0.0015,~\bar{\eta}_{2_{i}}=0.0012,~\bar{\varphi}_{1_{i}}=0.00001,~\bar{\varphi}_{2_{i}}=0.00001,~\gamma_{1_{i}}=0.01,~\gamma_{2_{i}}=0.99$. According to Theorem 5.4, as $\gamma_{1_{i}}=0.01>0$ and $0<\gamma_{2_{i}}=0.99<1$, $\mathcal{B}_{i}(x_{i},\vartheta_{i}):=\bar{\mathcal{B}}_{i}(x_{i})$ is an A-CSBC for each A-SHS with $\mu_{i}=0.049$, $\beta_{i}=5.0557$, $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau},\gamma_{2_{i}}\}=0.999$, $\varphi_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}=0.00001$ and $\eta_{i}\!=\!\max\{e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}=0.0012$, where condition (5.3) is fulfilled by selecting $\alpha_{1_{i}}=0.1,\alpha_{2_{i}}=8$. As per Theorem 4.2, $\mathcal{B}(x,\vartheta)=\sum_{i=1}^{1000}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})$ is the obtained A-CBC, satisfying the conditions in (3.2) with $\beta=\sum_{i=1}^{1000}\xi_{i}\beta_{i}=5.0557$, $\mu=\sum_{i=1}^{1000}\xi_{i}\mu_{i}=0.049$, $\eta=\sum_{i=1}^{N}\xi_{i}\eta_{i}=0.0012,$ and $\gamma=1+\Gamma=0.9929$, where $\xi_{i}\Gamma_{i}=-7.141\times\!10^{-6}\!<\!0$ satisfies condition (4.1b) with $\xi_{i}=0.001$$\forall i\in\{1,\dots,1000\}$.
Figure 2: Control barrier certificate (top) and closed-loop state trajectories (bottom) of a representative subsystem within the interconnected SHS $\Phi$ with 3-in-3-out topology with $10$ noise realizations for $a_{1}=-0.4,b_{1}=0.5,a_{2}=0.01,b_{2}=0.06$ (left), $a_{1}=-0.3,b_{1}=0.2,a_{2}=1.01,b_{2}=1$ (middle), $a_{1}=0.01,b_{1}=0.7,a_{2}=0.075,b_{2}=0.8$ (right). Initial and unsafe regions are depicted by green and red boxes. Moreover, $\bar{\mathcal{B}}_{i}(x_{i})=\bar{\mu}_{i}$ and $\bar{\mathcal{B}}_{i}(x_{i})=\bar{\beta}_{i}$ indicated by  and  , respectively.
For $a_{1}=-0.3$, $b_{1}=0.2$, $a_{2}=1.01$, and $b_{2}=1$, we construct a CBC and safety controller as $\bar{\mathcal{B}}_{i}(x_{i})=0.0008x^{4}_{i}-0.0074x^{3}_{i}+0.0222x^{2}_{i}-0.0284x_{i}+0.0195$ and $\nu_{i}=-1.02x^{3}_{i}+3.05$. The parameters satisfying (5.1a)–(5.1e) in Definition 5.1 are chosen as $\bar{\mu}_{i}=0.0199$, $\bar{\beta}_{i}=2.0147$, $\bar{\kappa}_{i}=7.0587\times 10^{-4}$, $\bar{\eta}_{1_{i}}=0.0028$, $\bar{\eta}_{2_{i}}=0.003$, $\bar{\varphi}_{1_{i}}=5.8374\times 10^{-6}$, $\bar{\varphi}_{2_{i}}=2.0981\times 10^{-6}$, $\gamma_{1_{i}}=0.04547$, and $\gamma_{2_{i}}=1.00001$. By Theorem 5.4, since $\gamma_{1_{i}}>0$ and $\gamma_{2_{i}}\geq 1$, the function $\mathcal{B}_{i}(x_{i},\vartheta_{i}):=e^{4.57\times 10^{-4}\vartheta_{i}}\bar{\mathcal{B}}_{i}(x_{i})$ defines an A-CSBC with $\mu_{i}=0.0199$, $\beta_{i}=2.0212$, $\gamma_{i}=0.9996$, $\varphi_{i}=2.0981\times 10^{-6}$, and $\eta_{i}=0.003$, using $\alpha_{1_{i}}=0.1$, $\alpha_{2_{i}}=8$. From Theorem 4.2, the A-CBC $\mathcal{B}(x,\vartheta)=\sum_{i=1}^{1000}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})$ satisfies the conditions in (3.2) with $\beta=2.0212$, $\mu=0.0199$, $\eta=0.003$, and $\gamma=0.9995$, where $\xi_{i}\Gamma_{i}=-4.5836\times 10^{-7}<0$ fulfills condition (4.1b), with $\xi_{i}=0.001$, $\forall i\in\{1,\dots,1000\}$.
For $a_{1}=0.01$, $b_{1}=0.7$, $a_{2}=0.075$, and $b_{2}=0.8$, we synthesize a CBC and safety controller as $\bar{\mathcal{B}}_{i}(x_{i})=0.0029x^{4}_{i}-0.0187x^{3}_{i}+0.0444x^{2}_{i}-0.0419x_{i}+0.027$ and $\nu_{i}=-0.103x^{3}_{i}+1.2$. The parameters satisfying (5.1a)–(5.1e) in Definition 5.1 are $\bar{\mu}_{i}=0.0978$, $\bar{\beta}_{i}=8.0719$, $\bar{\kappa}_{i}=0.0034$, $\bar{\eta}_{1_{i}}=0.01$, $\bar{\eta}_{2_{i}}=0.01$, $\bar{\varphi}_{1_{i}}=3.04\times 10^{-5}$, $\bar{\varphi}_{2_{i}}=1.0112\times 10^{-5}$, $\gamma_{1_{i}}=-0.31$, and $\gamma_{2_{i}}=0.77$. As $\gamma_{1_{i}}\leq 0$ and $0<\gamma_{2_{i}}<1$, Theorem 5.4 guarantees that $\mathcal{B}_{i}(x_{i},\vartheta_{i}):=0.77^{\frac{\vartheta_{i}}{8}}\bar{\mathcal{B}}_{i}(x_{i})$ defines an A-CSBC with $\mu_{i}=0.0778$, $\beta_{i}=7.8124$, $\gamma_{i}=0.9983$, $\varphi_{i}=1.0112\times 10^{-5}$, and $\eta_{i}=0.01$, using $\alpha_{1_{i}}=0.1$ and $\alpha_{2_{i}}=8$ to meet (5.3). By Theorem 4.2, the A-CBC $\mathcal{B}(x,\vartheta)=\sum_{i=1}^{1000}\xi_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})$ satisfies the conditions in (3.2) with $\beta=7.8124$, $\mu=0.0778$, $\eta=0.01$, and $\gamma=0.999$, where $\xi_{i}\Gamma_{i}=-9.7522\times 10^{-7}<0$ satisfies condition (4.1b) with $\xi_{i}=0.001$, $\forall i\in\{1,\dots,1000\}$.
By Theorem 3.4, the state trajectory of $\Phi$, initialized within $X_{0}=[0~2]^{1000}$, remains in the safe set $0~9]^{1000}$ over the time horizon $\mathcal{T}=100$ with probability at least $0.967$ (first bound), $0.8454$ (second bound), and $0.8710$ (third bound). A CBC for a representative subsystem of the interconnected SHS $\Phi$ is shown in Fig. 2. Closed-loop trajectories under different bounds with 10 noise realizations are presented in Fig. 2, where only discrete time steps are shown, consistent with the guaranteed safety intervals. The CBC and corresponding controllers for each bound were computed in approximately $11.64$ seconds using $66.38~\mathsf{Mbit}$ of memory on a MacBook Pro with an M2 Max chip and 32 GB RAM.
7.3 Comparison Study
Comparing the results of our approach applied to a loosely coupled network, such as a ring topology, and an almost dense topology—e.g., a 3-in-3-out interconnection topology—it becomes evident that several trends emerge. First and foremost, satisfying the compositional conditions becomes more challenging in all three cases as the value of $\xi_{i}\Gamma_{i}$ increases, i.e., it approaches zero. This arises from the fact that an increase in the dimensions of the adversarial input $w_{i}$ pushes the matrix $(-\Theta+\Psi)$ closer to the positive semidefinite cone.
Second, the safety probability associated with each bound decreases in the 3-in-3-out interconnection topology compared to its counterpart in the ring topology (with the same certificate template, i.e., a CBC of degree four). Ensuring the satisfaction of the safety property by tuning parameters within this fixed template comes at the cost of reduced probabilities. Higher safety probabilities could potentially be achieved by employing a more expressive template with additional coefficients; however, this would again lead to increased computational complexity.
Last but not least, both computational time and memory usage are higher for the denser interconnection topology, owing to the fact that, according to Lemma 6.1, the number of coefficients in the SOS polynomials $l_{w_{i}}(x_{i},\nu_{i},w_{i})$ and $\hat{l}_{w_{i}}(x_{i},\nu_{i},w_{i})$, which serve as Lagrangian multipliers, increases as higher connectivity directly leads to a higher dimensional adversarial input $w_{i}$, thereby amplifying the computational burden at the subsystem level.
7.4 Computational Complexity Analysis
In this subsection, we aim to thoroughly analyze the computational complexity of our proposed approach compared with the existing method in the relevant literature (Lavaei, Soudjani & Frazzoli, 2022).
By employing a compositional method, we reduce the computational complexity from polynomial, as is in the monolithic setting of (Lavaei, Soudjani & Frazzoli, 2022), to linear with respect to the number of subsystems in our work. Specifically, when trying to solve a safety problem through designing a polynomial CBC using SOSTOOLS, for a fixed degree $d$ of the CBC (e.g., CBC of degree four, as in our case study) and an interconnected network of total size $s=n_{i}\times N$, the number of coefficients to be designed is $\binom{s+d}{s}$, a combinatorial count. Hence, the computational complexity grows polynomially with $s$ (and consequently $N$) for a fixed $d$. In other words, the complexity grows as $\mathcal{C}\propto\mathcal{O}(s^{d})$, where $\mathcal{C}$ denotes the computational complexity.
As a result, the approach in (Lavaei, Soudjani & Frazzoli, 2022) necessitates solving the problem mentioned in our case study section over a monolithic system of dimension $1000$ (corresponding to $1000$ one-dimensional subsystems), resulting in a computational cost $\mathcal{C}\propto\mathcal{O}(1000^{4})$, which is not tractable due to the polynomial curse of dimensionality. In contrast, our method decomposes the network into $1000$ one-dimensional subsystems, solves the problem individually for each, and subsequently lifts the safety guarantees over the entire network using a compositional reasoning. Even though the per-subsystem complexity still remains polynomial, i.e., $\bar{\mathcal{C}}\propto\mathcal{O}(n_{i}^{d})$, with $\bar{\mathcal{C}}$ representing the computational complexity for each subsystem, the overall complexity scales linearly with the number of subsystems, yielding a total cost of $\mathcal{C}\propto\mathcal{O}(N)=N\times\bar{\mathcal{C}}$, or $\mathcal{C}\propto 1000\bar{\mathcal{C}}$ in the same example, where $\bar{\mathcal{C}}\propto\mathcal{O}(1^{4})$.
Figure 3: Comparison of the computational complexity between the monolithic approach in (Lavaei, Soudjani & Frazzoli, 2022) and our proposed compositional approach.
Remark 7.1.
In principle, the computations can be performed in parallel. Under a parallel computing framework, the total computational time becomes independent of the number of subsystems and only depends on the number of states, assuming access to adequate computational resources. For instance, solving the safety problem for a network of 1000 subsystems in parallel would require either 1000 individual processing units (e.g., personal computers) or a single high-performance computing platform (e.g., a GPU-equipped server).
8 Conclusion
In this work, we proposed a compositional framework for the safety controller synthesis of interconnected stochastic hybrid systems with both continuous and discrete evolutions. In our proposed setting, we first introduced an augmented framework to reformulate each stochastic hybrid subsystem with continuous and discrete evolutions with a unified framework covering both behaviors. We then introduced notions of augmented control sub-barrier certificates, constructed for each subsystem, by employing which we constructed augmented control barrier certificates for interconnected systems together with their safety controllers under some small-gain compositional conditions. By employing the constructed augmented control barrier certificate, we finally computed a guaranteed probabilistic bound on the safety of the interconnected stochastic hybrid systems.